### v2

# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

Abstract.
Famous results state that the classical MacWilliams identities fail for the Lee metric, the homogeneous metric and for the subfield metric, apart from some trivial cases. In this paper we change the classical idea of enumerating the codewords of the same weight and choose a finer way of partitioning the code that still contains all the information of the weight enumerator of the code. The considered decomposition allows for MacWilliams-type identities which hold for any additive weight over a finite chain ring. For the specific cases of the homogeneous and the subfield metric we then define a coarser partition for which the MacWilliams-type identities still hold. This result shows that one can, in fact, relate the code and the dual code in terms of their weights, even for these metrics. Finally, we derive Linear Programming bounds stemming from the MacWilliams-type identities presented.
Key words and phrases:
Lee weight, homogeneous weight, subfield weight, MacWilliams identity, Codes over rings
1. Introduction
The MacWilliams identities have been a major breakthrough in understanding the theory of linear codes and mark one of the most celebrated results. In a nutshell, the MacWilliams identities state that the weight enumerator of a code is related to, and completely determines, the weight enumerator of the dual code. The original results have been proven in [31] by Jessie MacWilliams and several generalizations and reformulations, through association schemes, generating functions or graphs, have appeared since [15, 27, 28, 42, 26]. Other generalizations of the MacWilliams identities have been discussed in [39] and further extensions to different code families, like convolutional codes, quantum codes and others, followed [11, 20, 21, 9, 46]. The MacWilliams identities also allow to state a Linear Programming (LP) bound, giving the tightest known upper bound on the size of a code with prescribed minimum distance and length. For an overview we refer the interested reader to [16].
While this powerful original result restricts itself to classical codes over finite fields endowed with the Hamming metric, many works have considered different ambient spaces and different metrics, with various outcomes. Some works on MacWilliams identities investigate partitions of the ambient space. Indeed, it is well-known that if the partition of the ambient space is Fourier-reflexive, the MacWilliams-type identities for this partition exist [19]. This last approach is mainly used for codes over rings and group codes, e.g. in [11, 25, 48]. The MacWilliams identities were generalized in many different ways, for instance, in [42] Wood employed a similar technique as described in [2] to generalize them to codes over Frobenius rings endowed with the Hamming metric. In 1972, Delsarte generalized them even further to additive codes using the Hamming metric [14].
However, as soon as we change the metric, the situation becomes more complex. While the MacWilliams identities are still valid for codes over $\mathbb{Z}/{4}\mathbb{Z}$ equipped with the Lee metric [24], in 2015, Shi et al. [37] and [43] proved their failure over $\mathbb{Z}/{8}\mathbb{Z}$. Moreover, Abdelghany and Wood [1] showed the nonexistence of any version of the MacWilliams identities for Lee weight enumerators over $\mathbb{Z}/{m}\mathbb{Z}$ for any positive integer $m\geq 5$. A similar result holds for the homogeneous weight as well: apart from the trivial cases $\mathbb{Z}/{p}\mathbb{Z}$, for a prime $p$, where the homogeneous metric coincides with the Hamming metric, the homogeneous MacWilliams identities fail [19, 44]. In both the cases of Lee and homogeneous metric, as well as for the subfield metric [22], one of the main problems when considering the weight distribution is the lack of knowledge about the entries of codewords for a fixed weight.
Changing the partition of a code and, more precisely, moving away from enumerating codewords of the same weight, has already been suggested by MacWilliams herself for Lee-metric codes over finite fields [32]. Similarly, this idea has been supported for Lee-metric codes using association schemes in [6, 41] where the existence of MacWilliams-type identities have been shown. The study of partitioning the ambient space according to a pre-defined property dates back to Zinoviev and Ericson [47] in 1996. They studied additive codes and analyzed for which enumerators of the code and its dual MacWilliams-type identities are fulfilled, which laid the foundation of Fourier-reflexive partitions. In [19], Gluessing-Luerssen studied partitions of codes over Frobenius rings and showed that MacWilliams identities exist if the partition of the underlying ambient space is Fourier-reflexive. In a follow-up work [18], she proved that partitioning a code in the homogeneous metric into the codewords of the same homogeneous weight yields to a failure of the MacWilliams identities. Similarly, for codes in the sum-rank metric, [10] showed the failure of the classical MacWilliams identities and the existence of a MacWilliams-type identities using a finer partition.
We place this paper in this line of research and generalize the MacWilliams identities further. Instead of enumerating tuples (or vectors) of a given weight, we introduce a novel partition and show that it allows to state MacWilliams-type identities for any finite chain ring and any additive weight. Remarkably, the considered partition allows to retrieve the weight enumerator. While this result holds in general (as shown also in [19, 45]), we additionally put a greater focus on the three metrics for which the classical MacWilliams identities famously fail, namely the Lee metric, the homogeneous metric and the subfield metric. This not only shows that in this setting the weight enumerator of a code is related to (and completely determines) the weight enumerator of the dual, but also allows to state linear programming bounds for the three considered metrics.
The paper is organized as follows. Section 2 will serve to introduce and recap the main definitions and results needed throughout the paper. We start with the notion of a linear code over a finite chain ring endowed with an additive weight (where we specifically introduce the Hamming, Lee, homogeneous and subfield weight). We then recap the theory of group characters stating some of their main properties, such as the Schur orthogonality. The section is completed by introducing the MacWilliams identities in their original form together with the Krawtchouk coefficients. Moving away from enumerating the codewords of the same weight, in Section 3 we present our novel results on the existence of MacWilliams-type identities. As a first step we consider a finite chain ring endowed with any additive metric and partition the code into codewords of the same (Lee weight) decomposition which we refer to as the Lee partition of the code. We show that the Lee decomposition enumerator of a code is fully determined by the Lee decomposition enumerator of its dual code, i.e., MacWilliams-type identities exist for any additive weight over any finite chain ring, and, in particular, for codes in the Lee metric. Even though the result holds for any additive metric over a finite chain ring, for some metrics there are coarser and more natural partitions to choose. Thus, in a next step we introduce partitions for the homogeneous weight and the subfield weight, respectively, that are more suited and coarser and still allow for MacWilliams-type identities. To complete the study, we derive linear programming bounds for the Lee, homogeneous and subfield metric in Section 4. Finally, concluding remarks are given in Section 5.
2. Preliminaries
In this section we introduce the main definitions and basic results needed throughout this paper. We denote by $p$ a prime number. Let $\mathcal{R}$ be a finite chain ring with maximal ideal generated by $\gamma$, let $s$ be its nilpotency index and $q=p^{r}$ be the size of the residue field, for a positive integer $r$. While the main result holds for any finite chain ring, we often focus on integer residue rings. We denote by $p$ a prime number and for a positive integer $s$ we let $\mathbb{Z}/{p^{s}}\mathbb{Z}$ denote the finite integer residue ring, and we denote by $(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$ the set of units modulo $p^{s}$. Note that a finite integer residue ring $\mathbb{Z}/{p^{s}}\mathbb{Z}$ is a finite chain ring with $q=p$, $\gamma=p$ and nilpotency index $s$.
Given a positive integer $n$ and an $\ell$-tuple of nonnegative integers $k:=(k_{1},\ldots,k_{\ell})$ satisfying that $\sum_{i=1}^{\ell}k_{i}=n$, we denote by
(1) $\binom{n}{k}:=\binom{n}{k_{1},\ldots,k_{\ell}}=\frac{n!}{k_{1}!\ldots k_{\ell}!}$
the multinomial coefficient. Moreover, given a statement $A$, we denote by $\mathbbm{1}_{A}$ its indicator function, that is,
(2) $\mathbbm{1}_{A}=\begin{cases}1&A\text{ is true,}\\ 0&\text{otherwise.}\end{cases}$
Furthermore, for a positive integer $m$, we denote by $I_{m}$ the $m\times m$ identity matrix.
2.1. Codes and Additive Weights over Finite Chain Rings
In this section we introduce linear codes over finite chain rings and their main properties. In this paper, we exclusively focus on commutative chain rings in order not to distinguish between left and right ideals.
Definition 1.
Let $\mathcal{R}$ be a finite chain ring. An $\mathcal{R}$-linear code $\mathcal{C}$ of length $n$ is an $\mathcal{R}$-submodule of $\mathcal{R}^{n}$. The free module $\mathcal{R}^{n}$ is called the ambient space of $\mathcal{C}$ and the elements of $\mathcal{C}$ are called codewords.
The fundamental theorem of finite abelian groups implies that any $\mathcal{R}$-module (i.e., any $\mathcal{R}$-linear code $\mathcal{C}$) is isomorphic to the following direct sum of $\mathcal{R}$-modules
(3) $\mathcal{C}\cong(\mathcal{R}/\gamma^{s}\mathcal{R})^{k_{0}}\times(\mathcal{R}/\gamma^{s-1}\mathcal{R})^{k_{1}}\times\cdots\times(\mathcal{R}/\gamma\mathcal{R})^{k_{s-1}}.$
The unique $s$-tuple $(k_{0},k_{1},\ldots,k_{s-1})$ is called the subtype of $\mathcal{C}$ and $k_{0}$ is called its free rank.
A code $\mathcal{C}\subseteq\mathcal{R}^{n}$ can be represented by a generating set, that is a subset of codewords that generates $\mathcal{C}$ as an $\mathcal{R}$-submodule. We call a generating set a minimal generating set if it is minimal with respect to inclusion. The cardinality of a minimal generating set of $\mathcal{C}$ is called the rank $K$ and can be computed as $K=\sum_{i=0}^{s-1}k_{i}$. This allows for a matrix representation in terms of a minimal generating set.
Definition 2.
Given an $\mathcal{R}$-linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ of free rank $k_{0}$ and rank $K$, a matrix $G\in\mathcal{R}^{K\times n}$ whose rows form a generating set of $\mathcal{C}$ is called a generator matrix of the code. A matrix $H\in\mathcal{R}^{(n-k_{0})\times n}$ whose kernel coincides with $\mathcal{C}$ is called a parity-check matrix of $\mathcal{C}$.
Usually, it is convenient to consider a generator and parity-check matrix in their standard form defined in [36].
Proposition 3 ([36, Proposition 3.2]).
Let $\mathcal{R}$ be a finite chain ring and let $\mathcal{C}\subseteq\mathcal{R}^{n}$ be a linear code of subtype $(k_{0},\dots,k_{s-1})$. Then $\mathcal{C}$ is permutation equivalent to a code having a generator matrix in the following standard form
(4) $G=\begin{bmatrix}I_{k_{0}}&A_{0,1}&A_{0,2}&A_{0,3}&\dots&A_{0,s-1}&A_{0,s}\\ 0&\gamma I_{k_{1}}&\gamma A_{1,2}&\gamma A_{1,3}&\dots&\gamma A_{1,s-1}&\gamma A_{1,s}\\ 0&0&\gamma^{2}I_{k_{2}}&\gamma^{2}A_{2,3}&\dots&\gamma^{2}A_{2,s-1}&\gamma^{2}A_{2,s}\\ \vdots&\vdots&\vdots&\vdots&&\vdots&\vdots\\ 0&0&0&0&\dots&\gamma^{s-1}I_{k_{s-1}}&\gamma^{s-1}A_{s-1,s}\\ \end{bmatrix}\ ,$
where $A_{i,s}\in(\mathcal{R}/\gamma^{s-i}\mathcal{R})^{k_{i}\times(n-K)}$ and $A_{i,j}\in(\mathcal{R}/\gamma^{s-i}\mathcal{R})^{k_{i}\times k_{j}}$ for $j<s$. Moreover $\mathcal{C}$ is permutation equivalent to a code having a parity-check matrix in standard form
(5) $H=\begin{bmatrix}B_{0,0}&B_{0,1}&\dots&B_{0,s-1}&I_{n-K}\\ \gamma B_{1,0}&\gamma B_{1,1}&\dots&\gamma I_{k_{s-1}}&0\\ \vdots&\vdots&&\vdots&\vdots\\ \gamma^{s-1}B_{s-1,0}&\gamma^{s-1}I_{k_{1}}&\dots&0&0\\ \end{bmatrix},$
where $B_{0,j}\in(\mathcal{R}/\gamma^{s}\mathcal{R})^{(n-K)\times k_{j}},B_{i,j}\in(\mathcal{R}/\gamma^{s-i}\mathcal{R})^{k_{s-i}\times k_{j}},$ for $j>1.$
We endow the ambient space $\mathcal{R}^{n}$ with the standard inner product, i.e., for $x,y\in\mathcal{R}^{n}$ we set
(6) $\left\langle{x,y}\right\rangle:=xy^{\top}=\sum_{i=1}^{n}x_{i}y_{i}.$
Definition 4.
Given a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$, the dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ is defined as
(7) $\mathcal{C}^{\perp}:=\left\{{x\in\mathcal{R}^{n}\;|\;\left\langle{x,c}\right\rangle=0\text{ for every }c\in\mathcal{C}}\right\}.$
In ring-linear coding theory, the notion of duality is in general not well-defined. However, in [42], Wood proved that the dual of a code over a Frobenius ring, and hence over a finite chain ring, is well-defined. Thus, it holds that $(\mathcal{C}^{\perp})^{\perp}=\mathcal{C}$. It is straightforward to see that a parity-check matrix $H$ of $\mathcal{C}$ is a generator matrix of $\mathcal{C}^{\perp}$. Moreover, if $\mathcal{C}$ is of rank $K$ and subtype $(k_{0},\dots,k_{s-1})$ then $\mathcal{C}^{\perp}$ has rank $n-k_{0}$ and subtype $(n-K,k_{s-1},\dots,k_{1})$.
One of the main goals of coding theory is to correct errors that occur during transmission in a received message. A linear code is capable of correcting a certain amount of errors which are measured using a given metric induced by a weight function over the ring $\mathcal{R}$.
Definition 5.
Let $\mathcal{R}$ denote a finite chain ring. A weight over $\mathcal{R}$ is a function $\mathrm{wt}\colon\mathcal{R}\to\mathbb{Q}$ satisfying
  1. (1)
    $\mathrm{wt}(0)=0$ and $\mathrm{wt}(x)>0$ for all $x\neq 0$;
  2. (2)
    $\mathrm{wt}(x)=\mathrm{wt}(-x)$;
  3. (3)
    $\mathrm{wt}(x+y)\leq\mathrm{wt}(x)+\mathrm{wt}(y)$.
A weight over a finite chain ring $\mathcal{R}$ naturally induces a distance which is a function $\mathrm{d}\colon\mathcal{R}\times\mathcal{R}\to\mathbb{Q}$ defined by $\mathrm{d}(x,y):=\mathrm{wt}(x-y)$. The coordinate-wise extensions of the weight and distance function will also be denoted by $\mathrm{wt}$ and $\mathrm{d}$ which we call (additive) weight and (additive) distance, respectively. In particular, given $x\in\mathcal{R}^{n}$, we define
(8) $\mathrm{wt}(x):=\sum_{i=1}^{n}\mathrm{wt}(x_{i}).$
It is easy to see that such an additive distance over $\mathcal{R}$ satisfies the properties of a metric. We will now introduce the main additive weight functions which are of interest in the course of this paper. Each of the weights over $\mathcal{R}$ induce a distance function as described above.
The most prominent weight in coding theory is the Hamming weight.
Definition 6.
Let $\mathcal{R}$ be a finite chain ring. The Hamming weight of an $n$-tuple $x\in\mathcal{R}^{n}$ is given by the cardinality of the support of $x$, i.e.,
(9) $\operatorname{wt}_{\mathsf{H}}(x):=\left|\,{\left\{{i=1,\ldots,n\;|\;x_{i}\neq 0}\right\}}\,\right|.$
We now focus on the case where $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$ is a finite integer residue ring modulo $p^{s}$. Hence, we consider codes to be submodules of $(\mathbb{Z}/p^{s}\mathbb{Z})^{n}$. In this work, we are particularly interested in the Lee weight and homogeneous weight. The Lee weight, introduced in 1958 by Lee [30], is a generalization of the Hamming weight over the binary.
Definition 7.
The Lee weight of an element $a\in\mathbb{Z}/{p^{s}}\mathbb{Z}$ is given by
(10) $\operatorname{wt}_{\mathsf{L}}(a):=\min\left\{a,\left|\,{p^{s}-a}\,\right|\right\}.$
Similarly, we define the Lee weight of an $n$-tuple $x\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ additively by
(11) $\operatorname{wt}_{\mathsf{L}}(x):=\sum_{i=1}^{n}\operatorname{wt}_{\mathsf{L}}(x_{i}).$
The following upper and lower bounds hold trivially for any $a\in\mathbb{Z}/{p^{s}}\mathbb{Z}$ and $x\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$
(12) $0\leq\operatorname{wt}_{\mathsf{L}}(a)\leq\left\lfloor{p^{s}/2}\right\rfloor\quad\text{and}\quad\operatorname{wt}_{\mathsf{H}}(x)\leq\operatorname{wt}_{\mathsf{L}}(x)\leq\left\lfloor{p^{s}/2}\right\rfloor\operatorname{wt}_{\mathsf{H}}(x).$
In 1997, Constantinescu and Heise first introduced the homogeneous weight for integer residue rings [13], which was later extended to finite rings in [23] and [35]. For this we consider again a finite chain ring $\mathcal{R}$ with socle $\mathcal{S}.$ Recall that the socle $\mathcal{S}$ of a ring is defined as the sum of the minimal nonzero submodules. For example, for $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$ we have $\mathcal{S}=\langle p^{s-1}\rangle$. More generally, if the maximal ideal of $\mathcal{R}$ is $\langle\gamma\rangle$, we have $\mathcal{S}=\langle\gamma^{s-1}\rangle.$
The homogeneous weight extends the Hamming weight coinciding with it for finite fields and with the Lee weight for $\mathbb{Z}/{4}\mathbb{Z}$. In the following we define the homogeneous weight in its normalized form.
Definition 8.
Let $\mathcal{R}$ be a finite chain ring with residue field of size $q$. The homogeneous weight of $a\in\mathcal{R}$ is defined as
(13) $\operatorname{wt}_{\mathsf{Hom}}(a)=\begin{cases}0&\text{if }a=0,\\ 1&\text{if }a\not\in\mathcal{S},\\ \frac{q}{q-1}&\text{if }a\in\mathcal{S}\setminus\{0\}.\end{cases}$
The homogeneous weight extends additively to tuples $x\in\mathcal{R}^{n}$, i.e.,
(14) $\operatorname{wt}_{\mathsf{Hom}}(x)=\sum_{i=1}^{n}\operatorname{wt}_{\mathsf{Hom}}(x_{i}).$
For the third metric, let us consider finite fields, more precisely finite extension fields $\mathbb{F}_{p^{r}}$. In [22], the authors introduced the subfield metric over $\mathbb{F}_{p^{r}}$ for quantum error correcting codes. This metric is defined similarly to the homogeneous metric over finite chain rings. Namely, it gives weight $1$ to nonzero elements in the underlying base field and all other nonzero elements in $\mathbb{F}_{p^{r}}$ are of weight $\lambda\geq 1$.
Definition 9.
Let $a\in\mathbb{F}_{p^{r}}$ and $\lambda\geq 1$. The $\lambda$-subfield weight of $a$ is defined as
(15) $\operatorname{wt}_{\lambda}(a)=\begin{cases}0&\text{if }a=0,\\ 1&\text{if }a\in\mathbb{F}_{p}^{\times},\\ \lambda&\text{if }a\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}.\end{cases}$
The $\lambda$-subfield weight extends additively to vectors $x\in\mathbb{F}_{p^{r}}^{n}$, i.e.,
(16) $\operatorname{wt}_{\lambda}(x)=\sum_{i=1}^{n}\operatorname{wt}_{\lambda}(x_{i}).$
2.2. Group Characters
We now recall the required basics on group characters. In a nutshell, a group character defines a representation of a group $G$ in terms of a complex function. Working over cyclic groups, characters are used in combination with the discrete Fourier transform.
Definition 10.
Let $G$ be a finite abelian group. A character $\chi$ of $G$ is a complex-valued map $\chi:G\longrightarrow\left\{{z\in\mathbb{C}\;|\;\left|\,{z}\,\right|=1}\right\}$, such that for every $\alpha,\beta\in G$ it holds
(17) $\chi(\alpha+\beta)=\chi(\alpha)\chi(\beta).$
A character is called principal if it maps every element of the group to $1$.
Example 11.
In the following, we derive explicit expressions for characters over certain finite abelian groups.
  1. (1)
    If $G=(\mathbb{Z}/{m}\mathbb{Z},+)$ with $m$ a positive integer, for any $a,x\in\mathbb{Z}/{m}\mathbb{Z}$ the characters of $G$ are given by
    (18) $\chi_{a}(x):=\xi^{ax}\ ,$
    where $\xi=e^{\frac{2\pi i}{m}}$ is an $m$-th root of unity.
  2. (2)
    Let $G$ be a finite abelian group. From the fundamental theorem of finite abelian groups, $G$ decomposes uniquely (up to the order of summands) into a finite direct sum of cyclic subgroups of prime power order, namely,
    (19) $G=\mathbb{Z}/{q_{1}}\mathbb{Z}\times\dots\times\mathbb{Z}/{q_{t}}\mathbb{Z}\ .$
    For any $a=(a_{1},\dots,a_{t}),\ x=(x_{1},\dots,x_{t})\in G$ any character has the form
    (20) $\chi_{a}(x)=\prod_{i=1}^{t}\xi_{i}^{a_{i}x_{i}}\ ,$
    where $\xi_{i}$ is a $q_{i}$-th root of unity. In particular, if $G=(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and $a,x\in G$ the characters are given by
    (21) $\chi_{a}(x)=\xi^{\left\langle{a,x}\right\rangle}\ ,$
    where $\left\langle{a,x}\right\rangle:=ax^{\top}$ and $\xi$ is a $p^{s}$-root of unity.
  3. (3)
    Let $G=(\mathbb{F}_{p^{r}}^{n},+)$ and $f:\mathbb{F}_{p^{r}}\longrightarrow\mathbb{F}_{p}$ a nontrivial $\mathbb{F}_{p}$-linear map. Any character of $G$ is given by $\chi_{a}(x)=\xi^{f(\left\langle{x,a}\right\rangle)}$, where $\xi$ is a primitive $p$-th root of unity and for $a,x\in G$, $\left\langle{a,x}\right\rangle=ax^{\top}$ denotes the standard inner product in $\mathbb{F}_{p^{r}}^{n}$. As an example of such non-trivial linear mapping, we can consider the expansion of an element $a\in\mathbb{F}_{p^{r}}$ to a vector $\Gamma(a)\in\mathbb{F}_{p}^{r}$ using a basis $\Gamma$ of $\mathbb{F}_{p^{r}}$ over $\mathbb{F}_{p}$ and project on the first coordinate of $\Gamma(a).$ Another example is the trace function $\operatorname{Tr}(x)=\sum_{i=0}^{m-1}x^{p^{i}},$ which we will use throughout this paper as $f$. Therefore, any character of $G$ has the form
    (22) $\chi_{a}(x)=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}\ ,$
    where $\xi$ is a primitive $p$-th root of unity.
  4. (4)
    Let $\mathcal{R}=\mathrm{GR}(p^{s},r)$ be the Galois ring of characteristic $p^{s}$ and with $p^{sr}$ elements, namely $\mathcal{R}$ is the Galois extension of $\mathbb{Z}/{p^{s}}\mathbb{Z}$ of degree $r$. Recall that any element $a\in\mathcal{R}$ has a unique $p$-adic representation, that is $a=a_{0}+a_{1}p+\dots+a_{s-1}p^{s-1}$, where, for any $11\leq i\leq s-$ 1, $a_{i}$ belongs to the Teichmüller set of $\mathcal{R}$ [33]. Over Galois rings, one can define the generalized Frobenius map as
    (23) $f\colon\mathcal{R}\to\mathcal{R},\quad a\mapsto a^{f}\coloneqq a_{0}^{p}+pa_{1}^{p}+\dots+p^{s-1}a_{s-1}^{p}\ .$
    Note that, if $r=1$, $f$ reduces to the Frobenius automorphism. The generalized Frobenius induces the generalized trace map
    (24) $\operatorname{Tr}\colon\mathcal{R}\to\mathbb{Z}/{p^{s}}\mathbb{Z},\quad a\mapsto a+a^{f}+a^{f^{2}}+\dots+a^{f^{r-1}}\ .$
    The generalized trace shares some properties with the trace map; for instance, the generalized trace is $\mathbb{Z}/{p^{s}}\mathbb{Z}$-linear, and it takes each value of $\mathbb{Z}/{p^{s}}\mathbb{Z}$ with equal frequency. For further details on this topic, refer to [40]. Similarly to the finite fields case, any character over $\mathcal{R}$ is of the form
    (25) $\chi_{a}(x)=\xi^{\operatorname{Tr}(ax)}\ ,$
    where $\xi$ is a $p^{s}$-th root of unity. Note that this generalizes the case of $\mathbb{Z}/{p^{s}}\mathbb{Z}$ when $r=1$, where the trace function coincides with the identity, as well as the case of $\mathbb{F}_{p^{r}}$ where $s=1$. In analogy with the finite abelian group case, characters in $\mathcal{R}^{n}$ can be explicitly described. Specifically, for any $a,x\in\mathcal{R}^{n}$ every character can be expressed as
    (26) $\chi_{a}(x)=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}\ ,$
    where $\xi$ is a $p^{s}$-th root of unity.
  5. (5)
    Let $\mathcal{R}$ be a finite chain ring of characteristic $p^{s}$. It is well-known that there exists a subring $\mathcal{T}\subseteq\mathcal{R}$ such that $\mathcal{R}$ is isomorphic to a direct product of multiple copies of $\mathcal{T}$ as abelian groups [12]. Therefore $\mathcal{R}=\prod_{i=1}^{t}\mathcal{T}$, where $\mathcal{T}$ is in particular a Galois ring and it is unique up to isomorphism. Given $\xi$ a $p^{s}$-th root of unity, for any $a=(a_{1},\dots,a_{t}),\ x=(x_{1},\dots,x_{t})\in\mathcal{R}$, a character $\chi_{a}(x)$ of $\mathcal{R}$ has the form
    (27) $\chi_{a}(x)=\prod_{i=1}^{t}\xi^{\operatorname{Tr}(a_{i}x_{i})}=\xi^{\operatorname{Tr}(a_{1}x_{1}+\dots+a_{t}x_{t})}=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}\ .$
    As in the previous cases, we can derive expressions for characters in $\mathcal{R}^{n}$. Given $a=(a^{(1)},\ldots,a^{(n)}),x=(x^{(1)},\ldots,x^{(n)})\in\mathcal{R}^{n},$ where $a^{(i)}=(a^{(i)}_{1},\ldots,a^{(i)}_{t})\in\mathcal{R}$ all the characters are of the form
    (28) $\chi_{a}(x)=\prod_{i=1}^{n}\xi^{\operatorname{Tr}\left({\left\langle{a^{(i)},x^{(i)}}\right\rangle}\right)}\ ,$
    where $\xi$ is a $p^{s}$-th root of unity.
An important property of characters is the Schur orthogonality. Given a character $\chi$ of a group $G$, the Schur orthogonality gives an explicit formula for the sum of the evaluations of $\chi$ in every element of the group $G$.
Lemma 12 (Schur Orthogonality over a Group).
Let $G$ be a finite group and $\chi$ a character of $G$. Then
(29) $\sum_{g\in G}\chi(g)=\begin{cases}\left|\,{G}\,\right|&\text{ if }\chi\text{ is the principal character,}\\ 0&\text{ otherwise}.\end{cases}$
Proof.
If $\chi$ is a principal character, we have $\chi(g)\equiv 1$ for every $g\in G$ and therefore $\sum_{g\in G}\chi(g)=\left|\,{G}\,\right|$. Thus, let us assume that $\chi$ is a non-principal character of $G$. This means that there exists an element $h\in G$ such that $\chi(h)\neq 1$. Since $G$ is a group, we have
(30) $\sum_{g\in G}\chi(g)=\sum_{g\in G}\chi(g+h)=\sum_{g\in G}\chi(g)\chi(h)=\chi(h)\sum_{g\in G}\chi(g),$
which is equivalent to $(1-\chi(h))\sum_{g\in G}\chi(g)=0$. As $\chi(h)\neq 1$ it must hold that $\sum_{g\in G}\chi(g)=0$. ∎
A similar result exists for subgroups (or linear codes). However, in that case, we distinguish between characters whose kernel is contained in the given subgroup and those whose kernel is not (see for instance [42]). More formally, let $H\subset G$ be a subgroup of a finite abelian group $G$. We denote by $\widehat{G}$ the set of all characters in $G$. Define the annihilator of $H\subset G$ as
(31) $(\widehat{G}:H):=\left\{{\chi\in\widehat{G}\;|\;\chi(h)=1\text{ for all }h\in H}\right\}.$
A character $\chi\in(\widehat{G}:H)$ is called a generating character. Then, the orthogonality relation in Lemma 12 extends to a subgroup as follows.
Lemma 13 (Schur Orthogonality over a Subgroup).
Let $G$ be a finite abelian group and $H\subset G$ be a subgroup. Then the following Schur orthogonality property holds
(32) $\sum_{h\in H}\chi(h)=\begin{cases}\left|\,{H}\,\right|&\text{ if }\chi\text{ is a generating character,}\\ 0&\text{ otherwise.}\end{cases}$
As a consequence of Schur’s Lemma 12, we highlight two relevant properties of characters that will be extensively used to prove MacWilliams-type identities.
Corollary 14.
Let $G$ be a finite abelian group and let $H$ be a subgroup of $G$. For a non-generating character $\chi$ (i.e., $\chi\not\in(\widehat{G}:H)$) it holds that
  • $\sum_{x\in H}\chi(x)=0$;
  • $\sum_{x\in H\setminus\{0\}}\chi(x)=-1$.
In particular, if $\mathcal{R}$ is a finite chain ring and $I$ is an ideal in $\mathcal{R}$, for a non-generating character $\chi$ over $\mathcal{R}$ we have $\sum_{x\in I}\chi(x)=0$, while $\sum_{x\in I\setminus\{0\}}\chi(x)=-1.$
We now turn our attention to partitions of abelian groups. Following [48], a partition $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ of $G$ consists of non-empty pairwise disjoint sets $P_{i}$, $1\leq i\leq m$, that cover $G$. The sets of a given partition are called blocks.
Definition 15.
Let $\mathcal{P}$ and $\mathcal{Q}$ be two partitions of $G$. The partition $\mathcal{P}$ is called finer than $\mathcal{Q}$ if the blocks of $\mathcal{P}$ are contained in the blocks of $\mathcal{Q}$. We also say that $\mathcal{Q}$ is coarser than $\mathcal{P}$ and we write $\mathcal{P}\leq\mathcal{Q}$.
We now define a partition of the set of characters $\widehat{G}$ of a finite abelian group which we refer to as the dual partition.
Definition 16.
Let $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ be a partition of a finite abelian group $G$. Let $\widehat{\mathcal{P}}=Q_{1}\mid Q_{2}\mid\dots\mid Q_{\ell}$ denote the partition of $\widehat{G}$ where $\chi,\chi^{\prime}\in\widehat{G}$ lie in the same block $Q_{j}$, for $j=1,\ldots,\ell$, if and only if, for every $i=1,\ldots,m$, it holds
(33) $\sum_{x\in P_{i}}\chi(x)=\sum_{x\in P_{i}}\chi^{\prime}(x).$
We call $\widehat{\mathcal{P}}$ the dual partition of $G$. Furthermore, a partition $\mathcal{P}$ is called self-dual if $\mathcal{P}=\widehat{\mathcal{P}}$.
In [19], Gluesing-Luerssen referred to a partition $\mathcal{P}$ as Fourier-reflexive if $\widehat{\widehat{\mathcal{P}}}=\mathcal{P}$, and showed that a partition $\mathcal{P}$ is Fourier-reflexive if and only if $\mathcal{P}$ and $\widehat{\mathcal{P}}$ consist of the same number of blocks. This definition aligns with the Fourier-invariant partitions introduced in [47].
In the following, we focus on self-dual partitions which, by definition, are Fourier-reflexive.
Definition 17.
Let $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ be a self-dual partition of a finite abelian group $G$. For $a\in P_{j}$, let $\chi_{a}$ be a character of $G$. If the sum $\sum_{x\in P_{i}}\chi_{a}(x)$ does not depend on the choice of $a\in P_{j}$, we define the Krawtchouk coefficient of $i$ with respect to $j$ as
(34) $K_{j}(i)=\sum_{x\in P_{i}}\chi_{a}(x)\ \text{ for any }a\in P_{j}\ .$
The existence of the Krawtchouk coefficients depends on the partition of the group, and [47, Lemma 1] shows that the Krawtchouk coefficients are well-defined if and only if the partitions are Fourier-invariant. Additionally, Gluesing-Luerssen in [19] showed that reflexive partitions lead to the MacWilliams identities.
In coding theory, partitions are usually induced by a weight function on the ambient space. Given a finite chain ring $\mathcal{R}$ and an additive weight $\operatorname{wt}$ on $\mathcal{R}^{n}$, we can partition $\mathcal{R}^{n}$ into tuples of the same weight. That is, we consider a partition $\mathcal{P}=P_{0}\;|\;\ldots\;|\;P_{\left|\,{\operatorname{Im}(\operatorname{wt})}\,\right|}$ where
(35) $P_{i}\coloneqq\{x\in\mathcal{R}^{n}\;|\;\operatorname{wt}(x)=i\}.$
Example 18.
Let $\mathcal{P}=P_{0}|\ldots|P_{n}$ be the Hamming weight partition of $\mathbb{F}_{q}^{n}$. That is, for any $0\leq i\leq n$,
(36) $P_{i}=\{x\in\mathbb{F}_{q}^{n}\;|\;\operatorname{wt}_{\mathsf{H}}(x)=i\}\ .$
Given $a\in P_{j}$, let $\chi_{a}$ be a character over $\mathbb{F}_{q}^{n}$. The Krawtchouk coefficients of the Hamming partition are well-defined and are described by the following explicit formula [16, Section 4.1],
(37) $K_{j}(i)=\sum_{x\in P_{i}}\chi_{a}(x)=\sum_{k=0}^{j}(-1)^{k}(q-1)^{j-k}\binom{i}{k}\binom{n-i}{j-k}\ .$
Therefore, the Hamming partition of $\mathbb{F}_{q}^{n}$ is Fourier-invariant.
2.3. The MacWilliams Identities over Finite Chain Rings
We finally recap the MacWilliams identities for the Hamming metric and the non-existence results in the Lee, homogeneous and subfield metric in this section. Again, let $\mathcal{R}$ be a finite chain ring and let $\mathrm{wt}\colon\mathcal{R}\longrightarrow\mathbb{Q}$ be an additive weight over $\mathcal{R}$. Furthermore, we consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$. From now on, we will denote by $M_{\mathrm{wt}}$ the maximum value of $\mathrm{wt}$ on $\mathcal{R}$.
The $\mathrm{wt}$-weight distribution of $\mathcal{C}$ is a tuple which specifies the number of codewords of each possible weight $i\in\left\{{0,\ldots,nM_{\operatorname{wt}}}\right\}$, i.e.,
(38) $A^{\operatorname{wt}}(\mathcal{C}):=\big{(}A_{0}^{\mathrm{wt}}(\mathcal{C}),A_{1}^{\mathrm{wt}}(\mathcal{C}),\dots,A_{nM_{\mathrm{wt}}}^{\mathrm{wt}}(\mathcal{C})\big{)},$
where $A_{i}^{\mathrm{wt}}(\mathcal{C}):=\left|\,{\left\{{c\in\mathcal{C}\;|\;\operatorname{wt}(c)=i}\right\}}\,\right|$. If $\mathrm{wt}$ and $\mathcal{C}$ are clear from the context, we will simply write $A_{i}$ instead of $A_{i}^{\mathrm{wt}}(\mathcal{C})$. Moreover, the weight distribution determines the $\mathrm{wt}$-weight enumerator of $\mathcal{C}$:
(39) $\mathsf{wwe}_{\mathcal{C}}=\sum_{i=0}^{nM_{\mathrm{wt}}}A_{i}^{\mathrm{wt}}(\mathcal{C})X^{nM_{\mathrm{wt}}-i}Y^{i}\ .$
That is, for instance, the Lee weight enumerator of a code $\mathcal{C}$ is denoted by $\mathsf{Lwe}_{\mathcal{C}}$, the homogeneous weight enumerator by $\mathsf{Homwe}_{\mathcal{C}}$ and the subfield weight enumerator by $\lambda\mathsf{we}_{\mathcal{C}}$. Although the weight distribution does not entirely characterize a code, it provides significant information. The research on this topic dates back to MacWilliams’ work [31], when she proved the famous MacWilliams identities (40). The identities establish a relation between the Hamming weight enumerators of a linear code over a finite field and its dual.
Theorem 19 (MacWilliams Identities [31, Lemma 2.2]).
Let $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ be a linear code over $\mathbb{F}_{q}$ and let $\mathcal{C}^{\perp}$ be its dual code. Let $(A_{i})_{i=0,\dots,n}$ and $(B_{i})_{i=0,\dots,n}$ be the Hamming weight distributions of $\mathcal{C}$ and $\mathcal{C}^{\perp}$, respectively. For any $0\leq\nu\leq n$ it holds
(40) $\sum_{j=0}^{n-\nu}\binom{n-j}{\nu}A_{j}=q^{n-\nu}\sum_{j=0}^{\nu}\binom{n-j}{n-\nu}B_{j}\ .$
There are several techniques to prove the MacWilliams identities each resulting in a distinct set of equations, but equivalent to the MacWilliams identities. One of the most widely used methods involves the Poisson summation formula [2]. On the other hand, Delsarte proposed an interesting approach based on association schemes and Krawtchouk coefficients [14].
Theorem 20.
Let $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ be a linear code and let $(A_{i})_{i=0,\dots,n}$ and $(B_{i})_{i=0,\dots,n}$ be the Hamming weight distributions of $\mathcal{C}$ and $\mathcal{C}^{\perp}$, respectively. For any $0\leq j\leq n$,
(41) $B_{j}=\frac{1}{|\mathcal{C}|}\sum_{i=0}^{n}A_{i}K_{j}(i),$
where the Krawtchouk coefficients $K_{j}(i)$ are the positive integers given in (37).
Some works on MacWilliams identities investigate partitions of the ambient space. Indeed, it is well-known that if the partition of the ambient space is Fourier-reflexive, the MacWilliams-type identities for this partition exist [19].
For the partition of $(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ into tuples of given Lee weight, however, the MacWilliams identities fail, whenever $p^{s}\geq 5$.
Theorem 21 ([1, Theorem 1.1]).
For any $p^{s}\geq 5$ there exist linear codes $\mathcal{C}_{1},\mathcal{C}_{2}$ over $\mathbb{Z}/{p^{s}}\mathbb{Z}$ satisfying $\mathsf{Lwe}_{\mathcal{C}_{1}}=\mathsf{Lwe}_{\mathcal{C}_{2}}$ and $\mathsf{Lwe}_{\mathcal{C}^{\perp}_{1}}\neq\mathsf{Lwe}_{\mathcal{C}^{\perp}_{2}}$.
Note that Theorem 21 has been shown for any $\mathbb{Z}/{m}\mathbb{Z}$ with $m\geq 5$. A similar result holds for the homogeneous weight as well.
Theorem 22 ([44, Theorem 3.2]).
For any prime number $p$ and any positive integer $s\geq 2$ with $p^{s}>5$, there exist linear codes $\mathcal{C}_{1},\mathcal{C}_{2}$ over $\mathbb{Z}/{p^{s}}\mathbb{Z}$ satisfying that $\mathsf{Homwe}_{\mathcal{C}_{1}}=\mathsf{Homwe}_{\mathcal{C}_{2}}$ and $\mathsf{Homwe}_{\mathcal{C}^{\perp}_{1}}\neq\mathsf{Homwe}_{\mathcal{C}^{\perp}_{2}}$.
In both the cases of Lee and homogeneous metric, as well as for the subfield metric (as shown in [22]), the Krawtchouk coefficients, introduced in Definition 17, are not well-defined as $\sum_{x\in P_{i}}\chi_{a}(x)$ depends on the choice of $a\in P_{j}$. Therefore, we cannot use Theorem 20 to obtain the MacWilliams identities. This leads us to considering a different partition of the ambient space.
In the following, we partition the ambient space based on a natural decomposition stemming from the Lee metric that allows us to partition any ambient space. The considered partition is finer than the weight partition and enables to recover any weight enumerator defined over the ambient space.
We want to emphasize, that this technique was already known to MacWilliams: she used a similar approach to show MacWilliams-type identities for codes over finite fields endowed with the Lee metric.
3. MacWilliams-Type Identities
As shown in Theorems 21 and 22, the MacWilliams identities do not exist for the Lee weight, for the homogeneous or the subfield weight, apart from some trivial cases. As highlighted, the problem mainly lies in the fact that, for the respective weight partitions,
(42) $\sum_{x\in P_{j}}\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}$
depends on the choice of $a\in P_{i}$, leading to Krawtchouk coefficients which are not well-defined.
In this section, we present MacWilliams-type identities for additive weights. More precisely, we start by showing the identities with respect to a partition suitable for the Lee metric which we refer to as the Lee partition. This approach and the MacWilliams identities resulting from it have already been found by MacWilliams herself in [32] over a finite field where she partitioned the code into codewords of the same Lee decomposition. In her work, MacWilliams represents the finite field $\mathbb{F}_{p^{r}}$ as $\mathbb{F}_{p^{r}}=\left\{{\omega_{0}=0,\pm\omega_{1},\ldots,\pm\omega_{\left\lfloor{p^{r}/2}\right\rfloor}}\right\}$ and refers to the Lee decomposition of a vector $v\in\mathbb{F}_{p^{r}}$ as the tuple whose $i$-th coordinate indicates the number of entries of $v$ equal to either $\omega_{i}\in\mathbb{F}_{p^{r}}$ or $-\omega_{i}\in\mathbb{F}_{p^{r}}$. It should be pointed out that the Lee metric is not properly defined on a non-prime finite field $\mathbb{F}_{p^{r}}$. Hence, over a finite field, this decomposition does not immediately recover the Lee weight of $v\in\mathbb{F}_{p^{r}}$. In [6, 41] similar MacWilliams-type identities for the Lee decomposition have been derived using association schemes. Additionally, it was shown in [19] that partitioning the ambient space into all singletons yields a Fourier-reflexive partition for which the MacWilliams identities exist. In the same paper it was shown that partitions based on the orbits of a subgroup of the automorphism group of the ambient space are also Fourier-reflexive. In [34] the case of symmetrized weight enumerators is discussed. We show that the identities hold for any finite chain ring $\mathcal{R}$ and any additive weight over $\mathcal{R}$. In particular, the identities hold also for the homogeneous and subfield metric.
3.1. Lee partition for Additive Weights
In the following, we consider $\mathcal{R}$ to be any finite chain ring with residue field size $q=p^{r}$, nilpotency index $s$ and $\mathcal{C}\subseteq\mathcal{R}^{n}$ a linear code. Moreover, let $\operatorname{wt}$ be an additive weight over $\mathcal{R}$. In this section, we label the elements of $\mathcal{R}$ as follows:
(43) =$\mathcal{R}=\{\alpha_{0}=0,\ \alpha_{1},\ \dots,\alpha_{M},\ \alpha_{M+1},\dots\ \alpha_{q^{s}-1}\},$
where $M\coloneqq\left\lfloor{\frac{q^{s}}{2}}\right\rfloor$ and $\alpha_{q^{s}-i}$ represents the additive inverse of $\alpha_{i}$ with respect to addition, i.e., $\alpha_{q^{s}-i}=-\alpha_{i}$. We refer to $\alpha_{q^{s}-i}$ as the opposite of $\alpha_{i}$.
Let us introduce the Lee partition of $\mathcal{R}$ where the blocks are formed by an element and its opposite.
Definition 23.
Given a finite chain ring $\mathcal{R}$, for any $0\leq i\leq M$ we set $I^{\mathsf{L}}_{i}\coloneqq\{\alpha_{i},-\alpha_{i}\}$ to be the subset of $\mathcal{R}$ containing $\alpha_{i}$ and its opposite. We then define the Lee partition of $\mathcal{R}$ as
(44) $\mathcal{I}^{\mathsf{L}}=I^{\mathsf{L}}_{0}\;|\;I^{\mathsf{L}}_{1}\;|\;\ldots\;|\;I^{\mathsf{L}}_{M}\ .$
The Lee partition $\mathcal{I}^{\mathsf{L}}$ of $\mathcal{R}$ gives rise to a Lee decomposition of a tuple $x\in\mathcal{R}^{n}$ capturing the number of entries equal to ±$\pm\alpha_{i}$ , for all $0\leq i\leq M$.
Definition 24.
Given a finite chain ring $\mathcal{R}$, let $\mathcal{I}^{\mathsf{L}}=I^{\mathsf{L}}_{0}\;|\;\ldots\;|\;I^{\mathsf{L}}_{M}$ be the Lee partition of $\mathcal{R}$. We define the Lee decomposition of an $n$-tuple $x\in\mathcal{R}^{n}$ as the $(M+1)$-tuple $\pi^{\mathsf{L}}(x)=(\pi^{\mathsf{L}}_{0}(x),\ldots,\pi^{\mathsf{L}}_{M}(x))$, where for each $i\in\left\{{0,\ldots,M}\right\}$
(45) $\pi^{\mathsf{L}}_{i}(x)=\left|\,{\left\{{k=1,\ldots,n\;|\;x_{k}\in I^{\mathsf{L}}_{i}}\right\}}\,\right|.$
Therefore, the Lee partition of $\mathcal{R}$ induces a partition of the ambient space $\mathcal{R}^{n}$.
Definition 25.
Let $\mathcal{R}$ be a finite chain ring and $n$ a positive integer. The set of all Lee decompositions in $\mathcal{R}^{n}$ is denoted by
(46) $\mathbb{D}^{\mathsf{L}}_{p,r,s,n}:=\left\{{\pi\in\left\{{0,\ldots,n}\right\}^{M}\,\Big{|}\,\sum_{i=0}^{M}\pi_{i}=n}\right\}.$
Moreover, given $\mathbb{D}^{\mathsf{L}}_{p,r,s,n}=\left\{{\pi^{(0)},\dots,\pi^{(D-1)}}\right\}$, for some positive integer $D$, we define the Lee partition of $\mathcal{R}^{n}$ as
(47) $\mathcal{P}^{\mathsf{L}}=P^{\mathsf{L}}_{\pi^{(0)}}\;|\;P^{\mathsf{L}}_{\pi^{(1)}}\;|\;\ldots\;|\;P^{\mathsf{L}}_{\pi^{(D-1)}}.$
That is, for every $i\in\left\{{0,\ldots,D-1}\right\}$
(48) $P^{\mathsf{L}}_{\pi^{(i)}}\coloneqq\left\{{x\in\mathcal{R}^{n}\;|\;\pi^{\mathsf{L}}(x)=\pi^{(i)}}\right\}.$
Given two Lee decompositions $\pi,\rho\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, with the aim of computing the component-wise product of two tuples with Lee decompositions $\pi$ and $\rho$, respectively, we denote the set of all compositions of $\rho$ with respect to $\pi$ by
(49) $\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho):=\Big{\{}t=(t_{0},\ldots,t_{M})\,\Big{|}\,$ $t_{i}=(t_{i0},\ldots,t_{iM})\in\left\{{0,\ldots,\pi_{i}}\right\}^{M}\text{ with }$
(50) $\sum_{j=0}^{M}t_{ij}=\pi_{i}\text{ and }\sum_{i=0}^{M}t_{ij}=\rho_{j},\;\forall i,j\in\left\{{0,\ldots,M}\right\}\Big{\}}.$
Figure 1 illustrates the set $\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$, i.e., the compositions of $\rho\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$ with respect to $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$.
[…]
Figure 1. Illustration of $\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$.
Different to the original MacWilliams identities for the Hamming metric, we aim at enumerating all codewords of the same Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$. That is, for each $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, we are interested in
(51) $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})\coloneqq\left|\,{P^{\mathsf{L}}_{\pi}\cap\mathcal{C}}\,\right|.$
Furthermore, if $\mathbb{D}^{\mathsf{L}}_{p,r,s,n}=\left\{{\pi^{(0)},\dots,\pi^{(D-1)}}\right\}$, we denote by
(52) $\mathcal{D}^{\mathsf{L}}(\mathcal{C}):=\left(\mathcal{D}_{\pi^{(0)}}^{\mathsf{L}}(\mathcal{C}),\ldots,\mathcal{D}_{\pi^{(D-1)}}^{\mathsf{L}}(\mathcal{C})\right)$
the Lee decomposition enumerator. If the code $\mathcal{C}$ considered is clear from the context, we will simply write $\mathcal{D}^{\mathsf{L}}{}$ and $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}{}$, for $i\in\left\{{0,\ldots,D-1}\right\}$.
Note that the Lee partition is always finer than the partition induced by any additive weight $\operatorname{wt}$. This is because, for any weight $\operatorname{wt}$, the weight of each element must be equal to the weight of its opposite. Consequently, the blocks of the partition of $\mathcal{R}^{n}$ induced by $\operatorname{wt}$ contain the blocks of the Lee partition. Thus, for an $n$-tuple $x\in\mathcal{R}^{n}$ we can compute its weight $\operatorname{wt}(x)$ using its Lee decomposition $\pi^{\mathsf{L}}(x)$, i.e.,
(53) $\operatorname{wt}(x)=\sum_{i=0}^{M}\pi^{\mathsf{L}}_{i}(x)\operatorname{wt}(i).$
By abuse of notation, we call this the $\operatorname{wt}$-weight of $\pi$ and we write $\operatorname{wt}(\pi)$. Thus, given any code $\mathcal{C}\subseteq\mathcal{R}^{n}$ with weight enumerator $A^{\operatorname{wt}}(\mathcal{C})=(A^{\operatorname{wt}}_{0},\ldots,A^{\operatorname{wt}}_{nM_{\mathrm{wt}}})$ as defined in (38) we have
(54) $A_{i}^{\operatorname{wt}}(\mathcal{C})=\sum_{\begin{subarray}{c}\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}\\ \operatorname{wt}(\pi)=i\end{subarray}}\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C}).$
It is important to note that the Lee decomposition of a codeword gives us information about both the size of the codeword’s support and composition. Additionally, the Lee partition can be interpreted as the orbits of the action of the group {±$\left\{{\pm 1}\right\}$ on the ring $\mathcal{R}$. It has been shown in [19, Theorem 2.6 and Theorem 2.7] that a partition $\mathcal{P}_{H}$ of finite group $G$ defined by the orbits of a subgroup $H\leq\mathrm{Aut}(G)$ is reflexive and allows for a MacWilliams identity. This immediately implies that the MacWilliams-type identities for the Lee decompositions exist for every finite chain ring $\mathcal{R}$ endowed with any additive weight.
In particular, if $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$, i.e., $r=1$, for all $i\in\left\{{0,\dots,M}\right\}$, the sets $I_{i}^{\mathsf{L}}$ reduce to $I_{i}^{\mathsf{L}}=\left\{{\pm i}\right\}$, $\operatorname{Tr}=\mathrm{id}$, and [19, Theorem 2.7] provides MacWilliams-type identities for Lee-metric codes.
Corollary 26 (MacWilliams-type Identities in the Lee Metric).
For a positive integer $n$, consider a linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and its dual $\mathcal{C}^{\perp}$. For a given Lee decomposition $\rho\in\mathbb{D}_{p,1,s,n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of Lee decomposition $\rho$ is given by
(55) $\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C}),$
where $K^{\mathsf{L}}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$ and, for a given $p^{s}$-th root of unity $\xi$, it is given by
(56) $K^{\mathsf{L}}_{\pi}(\rho)\coloneqq\begin{cases}\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\left(\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M-1}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\xi^{iM}\right)&\text{if }p=2,\\ \sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\left(\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\right)&\text{otherwise.}\\ \end{cases}$
To illustrate the calculations of the identities, let us now give an example for the Lee-metric MacWilliams-type identities.
Example 27.
Consider the code $\mathcal{C}\subseteq(\mathbb{Z}/{9}\mathbb{Z})^{3}$ with generator matrix $G=\begin{pmatrix}3&2&8\end{pmatrix}$. That is,
[…]
Note that there are five distinct Lee decompositions in $\mathcal{C}$, namely,
(58) $\pi^{(0)}=(3,0,0,0,0),\ \pi^{(1)}=(1,0,0,2,0),\ \pi^{(2)}=(0,1,1,1,0),\ \pi^{(3)}=(0,0,1,1,1)\text{ and }\pi^{(4)}=(0,1,0,1,1).$
Additionally, we have that $\mathcal{D}_{\pi^{(0)}}^{\mathsf{L}}(\mathcal{C})=1$ and $\mathcal{D}_{\pi^{(j)}}^{\mathsf{L}}(\mathcal{C})=2$ for $j=1,2,3,4$. The dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ is generated by the matrix
(59) $G^{\perp}=\begin{pmatrix}1&0&3\\ 0&1&2\end{pmatrix}.$
Let $\rho=(2,0,0,1,0)$ be a Lee decomposition in $(\mathbb{Z}/{9}\mathbb{Z})^{3}$. By calculating the elements of $\mathcal{C}^{\perp}$, we observe two codewords in $\mathcal{C}^{\perp}$ of Lee decomposition $\rho$, namely $(3,0,0)$ and $(6,0,0)$. Hence, $\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C}^{\perp})=2$.
We now compute the Krawtchouk coefficient $K^{\mathsf{L}}_{\pi}(\rho)$, as stated in (56), for $\pi=\pi^{(j)}$ with $j=0,\ldots,4$. Let us start with $\pi=\pi^{(0)}=(3,0,0,0,0)$. Notice, since $\pi_{i}=0$ for every $i=1,\ldots,4$, the only possibility for $t=(t_{0},\ldots t_{M})$ is given by $t_{0}=\rho$ and $t_{i}=(0,\ldots,0)$. That is, $t_{ij}=0$ and thus $(\xi^{-ij}+\xi^{ij})^{t_{ij}}=1$ for every $i>0$ and every $j$, as well as $\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}=1$. Hence, for $\pi=\pi^{(1)}$ we have
(60) $\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}=\binom{\pi_{0}}{t_{00},\ldots,t_{0M}}\left(1+1\right)^{t_{03}}=\binom{3}{2,1}(1+1)=6.$
Consider now $\pi=\pi^{(1)}=(1,0,0,2,0)$. There are two compositions $t,s\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$, namely
(61) $t=(t_{0},\ldots,t_{M})$ $:\ t_{0}=(1,0,0,0,0),t_{3}=(1,0,0,1,0)\text{ and }t_{1}=t_{2}=t_{4}=(0,\ldots,0),$
(62) $s=(s_{0},\ldots,s_{M})$ $:\ s_{0}=(0,0,0,1,0),s_{3}=(2,0,0,0,0)\text{ and }s_{1}=s_{2}=s_{4}=(0,\ldots,0).$
Hence, for $\pi=\pi^{(1)}$ we get
(63) $\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}$ $\left(\prod_{i=0}^{4}\binom{\pi_{i}}{t_{i0},\ldots,t_{i4}}\prod_{j=1}^{4}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\right)$
(64) $=\left(\binom{\pi_{0}}{t_{00},\ldots,t_{04}}\prod_{j=1}^{4}\left(\xi^{0}+\xi^{0}\right)^{t_{0j}}\binom{\pi_{3}}{t_{30},\ldots,t_{34}}\prod_{j=1}^{M}\left(\xi^{-3j}+\xi^{3j}\right)^{t_{3j}}\right.$
(65) $\qquad+\left.\binom{\pi_{0}}{s_{00},\ldots,s_{04}}\prod_{j=1}^{4}\left(\xi^{0}+\xi^{0}\right)^{s_{0j}}\binom{\pi_{3}}{s_{30},\ldots,s_{34}}\prod_{j=1}^{M}\left(\xi^{-3j}+\xi^{3j}\right)^{s_{3j}}\right)$
(66) $=\left(\binom{1}{1,0,0,0,0}\binom{2}{1,0,0,1,0}(\xi^{0}+\xi^{0})^{1}\binom{1}{0,0,0,1,0}\binom{2}{2,0,0,0,0}(\xi^{0}+\xi^{0})^{1}\right)$
(67) $=6.$
For $\pi=\pi^{(2)},\pi^{(3)},\pi^{(4)}$ the computations are similar and yield in each case
(68) $\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\left(\prod_{i=0}^{4}\binom{\pi_{i}}{t_{i0},\ldots,t_{i4}}\prod_{j=1}^{4}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\right)=\left((\xi^{0}+\xi^{0})+(\xi^{-3}+\xi^{3})+(\xi^{-3}+\xi^{3})\right)=0,$
since $\sum_{i\in\left\langle{3}\right\rangle}\xi^{i}=0$.
Putting the expressions given in (60), (63) and (68) in the right-hand side of (55) yields
(69) $\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})=\frac{1}{9}\left(6\cdot 1+6\cdot 2+0\cdot 2+0\cdot 2+0\cdot 2\right)=2,$
which, indeed, corresponds to $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C}^{\perp})=2$.
Compared to other partitions (like, for instance, partitioning a code $\mathcal{C}\subseteq\mathcal{R}^{n}$ by the codewords of the same weight), the partition into codewords of the same Lee decomposition is rather fine. The finer a partition is, the more complex the enumerator and the identities become. For a linear Lee-metric code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$, however, the Lee partition $\mathcal{P}^{\mathsf{L}}=P^{\mathsf{L}}_{\pi^{(0)}}\;|\;\ldots\;|\;P^{\mathsf{L}}_{\pi^{(D-1)}}$ is the coarsest partition for which MacWilliams-type identities exist and which allows to derive the Lee weight enumerator of $\mathcal{C}$, since for every $c\in P_{\pi^{(i)}}$ we have that $\operatorname{wt}_{\mathsf{L}}(c)=\operatorname{wt}_{\mathsf{L}}(\pi^{(i)})$. In fact, any coarser partition $\mathcal{Q}\geq\mathcal{P}^{\mathsf{L}}$ of $\mathcal{C}$ for which the Krawtchouk coefficient is well-defined (and hence MacWilliams-type identities exist) does not give back the Lee weight enumerator of $\mathcal{C}$. A simple example of this would be the partition into the Hamming weights. These partitions are not of our interest, as the goal is to find MacWilliams-type identities for a partition of the code that is in relation with the weight enumerator of a code.
Even though the MacWilliams identities for the Lee partition are valid for any additive weight over a finite chain ring $\mathcal{R}$, it is not always the most practical partition to choose. Indeed, for other metrics, there exist coarser partitions of the ambient space that still allow for MacWilliams-type identities and allow us to recover the weight enumerator of the code. In the following sections, we focus on the homogeneous metric and the subfield metric, and we introduce, for both metrics respectively, a coarser (and, hence, more practical) partition for which MacWilliams identities exist.
3.2. The Homogeneous Metric
We now focus on the homogeneous metric over $\mathcal{R}$, as defined in Definition 8, with residue field size $q=p^{r}$ and nilpotency index $s>1$. In fact, when $s=1$, $\mathcal{R}$ is a finite field, and the homogeneous metric reduces to the Hamming metric, for which MacWilliams identities exist. On the other hand, if $s>1$, MacWilliams identities for the homogeneous weight enumerator in the classical sense do not exist [18, 44]. Similar to the Lee metric, $n$-tuples with a given homogeneous weight can be decomposed in different ways. A natural choice is the homogeneous decomposition, enumerating the number of entries of $x\in\mathcal{R}^{n}$ of given homogeneous weight. That is, given an $n$-tuple $x\in\mathcal{R}^{n}$, its homogeneous decomposition is defined to be a triple $(\pi^{\mathsf{Hom}}_{0}(x),\pi^{\mathsf{Hom}}_{1}(x),\pi^{\mathsf{Hom}}_{q/(q-1)})$ where for each $i\in\left\{{0,1,\frac{q}{q-1}}\right\}$ we set
(70) $\pi^{\mathsf{Hom}}_{i}(x):=\left|\,{\left\{{k=1,\ldots,n\;|\;\operatorname{wt}_{\mathsf{Hom}}(x_{k})=i}\right\}}\,\right|.$
Hence, we can define the homogeneous decomposition of a code $\mathcal{C}\subseteq\mathcal{R}^{n}$, partitioning $\mathcal{C}$ into the codewords of the same homogeneous decomposition. We denote the homogeneous decomposition enumerator of $\mathcal{C}$ by $\mathsf{Homdec}_{\mathcal{C}}$. However, this partition does not lead to well-defined Krawtchouk coefficients [18].
One main reason is that, given an element $a$ of homogeneous weight $1$ and a $p^{s}$-th root of unity $\xi$, the sum
(71) $\sum_{\begin{subarray}{c}x\in\mathcal{R}\\ \operatorname{wt}_{\mathsf{Hom}}(x)=1\end{subarray}}\xi^{\operatorname{Tr}(ax)}$
is not independent on $a$. In fact, its value depends on whether $a$ and $x$ are units or zero divisors. Therefore, we refine it by distinguishing between zero components, units, nonzero components lying in the socle $\mathcal{S}=\left\langle{\gamma^{s-1}}\right\rangle$ of $\mathcal{R}$, and components lying in $\left\langle{\gamma}\right\rangle$ but not in $\mathcal{S}$, i.e., we define the sets
(72) $Z:=\left\{{0}\right\},\;U:=\mathcal{R}^{\times},\;S:=\mathcal{S}\setminus\left\{{0}\right\},\;\text{and}\;R:=\mathcal{R}\setminus\left(Z\cup U\cup S\right)=\langle\gamma\rangle\setminus\langle\gamma^{s-1}\rangle.$
We observe that the cardinalities of the sets $Z,U,S$ and $R$, respectively, are given by
(73) $\left|\,{Z}\,\right|=1,\;\left|\,{U}\,\right|=q^{s-1}(q-1),\;\left|\,{S}\,\right|=q-1,\;\text{and}\;\left|\,{R}\,\right|=q^{s-1}-q.$
This immediately partitions $\mathcal{R}$ into these four sets. We call this partition of $\mathcal{R}$ the homogeneous weight-unit partition and denote it by
(74) $\mathcal{I}^{\mathsf{Hom}}:=Z\;|\;U\;|\;S\;|\;R.$
The homogeneous weight-unit partition $\mathcal{I}^{\mathsf{Hom}}$ of $\mathcal{R}$ allows to define the homogeneous weight-unit decomposition of a tuple $x\in\mathcal{R}^{n}$.
Definition 28.
Let $\mathcal{R}$ be a finite chain ring. We define the homogeneous weight-unit decomposition of an element $x\in\mathcal{R}^{n}$ as a tuple $\pi^{\mathsf{Hom}}(x):=(\pi^{\mathsf{Hom}}_{Z}(x),\pi^{\mathsf{Hom}}_{U}(x),\pi^{\mathsf{Hom}}_{S}(x),\pi^{\mathsf{Hom}}_{R}(x))$, where
(75) $\pi^{\mathsf{Hom}}_{I}(x)=\left|\,{\left\{{k=1,\ldots,n\;|\;x_{k}\in I}\right\}}\,\right|,\text{ for }I\in\left\{{Z,U,S,R}\right\}.$
We denote the set of all homogeneous weight-unit decompositions in $\mathcal{R}^{n}$ by
(76) $\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}=\left\{{\pi\in\left\{{0,\ldots,n}\right\}^{4}\Big{|}\sum_{I\in\left\{{Z,U,S,R}\right\}}\pi_{I}=n}\right\}.$
Definition 29.
If $\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}=\left\{{\pi^{(0)},\dots,\pi^{(D-1)}}\right\}$, for some positive integer $D$ is the set of all homogeneous weight-unit decompositions in $\mathcal{R}^{n}$, we define the homogeneous weight-unit partition of $\mathcal{R}^{n}$ as
(77) $\mathcal{P}^{\mathsf{Hom}}:=P^{\mathsf{Hom}}_{\pi^{(0)}}\;|\;P^{\mathsf{Hom}}_{\pi^{(1)}}\;|\;\ldots\;|\;P^{\mathsf{Hom}}_{\pi^{(D-1)}},$
where, for every $i\in\left\{{0,\ldots,D-1}\right\}$,
(78) $P^{\mathsf{Hom}}_{\pi^{(i)}}\coloneqq\left\{{x\in\mathcal{R}^{n}\;|\;\pi^{\mathsf{Hom}}(x)=\pi^{(i)}}\right\}.$
Finally, given two homogeneous weight-unit decompositions $\pi,\rho\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}$, we denote the set of all compositions of $\rho$ with respect to $\pi$ as
(79) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho):=\Big{\{}t=(t_{Z},t_{U},t_{S},t_{R})\ \big{|}$ $\ t_{i}=(t_{iZ},\ldots,t_{iR})\in\left\{{0,\ldots,\pi_{i}}\right\}^{4}\text{ with }$
(80) $\sum_{j\in\left\{{Z,U,S,R}\right\}}t_{ij}=\pi_{i}\text{ and }\sum_{i\in\left\{{Z,U,S,R}\right\}}t_{ij}=\rho_{j}\ \forall i,j\Big{\}}.$
Given an $\mathcal{R}$-linear code $\mathcal{C}$, the number of codewords $c\in\mathcal{C}$ of homogeneous weight-unit decomposition $\pi\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}$ is computed as
(81) $\mathcal{D}_{\pi}^{\mathsf{Hom}}(\mathcal{C}):=\left|\,{P^{\mathsf{Hom}}_{\pi}\cap\mathcal{C}}\,\right|.$
Furthermore, note that every $c\in P^{\mathsf{Hom}}_{\pi}$ has the same homogeneous weight given by $\operatorname{wt}_{\mathsf{Hom}}(c)=\pi_{U}+\pi_{R}+\frac{p}{p-1}\pi_{S}$. Hence, the homogeneous weight enumerator $\mathsf{Homwe}_{\mathcal{C}}$ can immediately be derived from the homogeneous weight-unit enumerator. MacWilliams-type identities exists for the homogeneous weight-unit enumerator.
Theorem 30 (MacWilliams-type Identities for the Homogeneous Weight-Unit Enumerator).
Let $\mathcal{R}$ be a finite chain ring with residue field size $q=p^{r}$ and nilpotency index $s$. Consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given homogeneous weight-unit decomposition $\rho\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of homogeneous weight-unit decomposition $\rho$ is given by
(82) $\mathcal{D}_{\rho}^{\mathsf{Hom}}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}}K^{\mathsf{Hom}}_{\pi}(\rho)\mathcal{D}_{\pi}^{\mathsf{Hom}}(\mathcal{C}).$
Here $K^{\mathsf{Hom}}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$, and for a given $p^{s}$-th root of unity $\xi$ is given by
(83) $K^{\mathsf{Hom}}_{\pi}(\rho)=\sum_{t\in\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho)}\Bigg{(}$ $\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}\Bigg{)}(-1)^{t_{US}}(-q^{s-1})^{t_{SU}}\left(q^{s-1}(q-1)\right)^{t_{ZU}}$
(84) $\left(q^{s-1}-q\right)^{t_{ZR}}(q-1)^{t_{ZS}+t_{SS}+t_{SR}+t_{RS}}(-q)^{t_{RR}}\mathbbm{1}_{\left\{{t_{UU}=t_{UR}=t_{RU}=0}\right\}}.$
Proof.
We start by proving that, for a given $a\in\mathcal{R}^{n}$ with homogeneous weight-unit decomposition $\pi$, the sum $\sum_{x\in P^{\mathsf{Hom}}_{\rho}}\chi_{a}(x)=\sum_{x\in P^{\mathsf{Hom}}_{\rho}}\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}$ provides the correct expression for the Krawtchouk coefficients, which are indeed well-defined. Note that the summation is made over all $x\in\mathcal{R}^{n}$ with given homogeneous weight-unit decomposition $\rho$. Furthermore, it is only dependent on the homogeneous weight-unit decomposition, but not on the composition of the tuples itself. Therefore, without loss of generality, we assume that the first $\pi_{Z}$ entries of $a$ have weight 00, the next $\pi_{U}$ entries are units. We then find $\pi_{S}$ entries lying in $S=\langle\gamma^{s-1}\rangle\setminus\{0\}$ and finally $\pi_{R}$ entries lying in $R=\left\langle{\gamma}\right\rangle\setminus\left\langle{\gamma^{s-1}}\right\rangle$.
In the first $\pi_{Z}$ positions, an $n$-tuple $x\in\mathcal{R}^{n}$ of homogeneous-weight unit decomposition $\rho$ can have $t_{Zj}<\rho_{j}$ positions lying in $j\in\left\{{Z,U,S,R}\right\}$, where $\sum_{j\in\left\{{Z,U,S,R}\right\}}t_{Zj}=\pi_{Z}$. Repeating the argument iteratively for each $i\in\left\{{Z,U,S,R}\right\}$, we deduce that the number of $x$ satisfying these conditions is then given by the product of multinomial coefficients
(85) $\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}.$
Now, for any position $k\in\left\{{1,\ldots,n}\right\}$, for each $I,J\in\left\{{Z,U,S,R}\right\}$ and $a_{k}\in I$, we compute
(86) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}.$
It is easy to see that if $J=Z$, for $I\in\left\{{Z,U,S,R}\right\}$ and any $a_{k}\in I$, we get
(87) $\sum_{x_{k}\in Z}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\xi^{0}=1.$
Furthermore, note that if $I=Z$ we have $a_{k}=0$. Thus, for every $J\in\left\{{Z,U,S,R}\right\}$ we obtain
(88) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\sum_{y\in J}\xi^{0}=\left|\,{J}\,\right|,$
where the cardinalities $\left|\,{J}\,\right|$ with $J\in\left\{{Z,U,S,R}\right\}$ are given in (73).
Similarly, for any $(I,J)\in\left\{{(S,S),(S,R),(R,S)}\right\}$ we have that the product $a_{k}x_{k}=0$ where $a_{k}\in I$ and $x_{k}\in J$. Therefore, we obtain again
(89) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\sum_{y\in J}\xi^{0}=\left|\,{J}\,\right|.$
Let us now consider $I=U$ the set of units and $a_{k}\in I$. For any $J\in\left\{{U,S,R}\right\}$ we obtain $a_{k}J=J$. Hence, applying Corollary 14, we get
(90) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\sum_{y\in J}\xi^{\operatorname{Tr}(y)}=\begin{cases}\sum_{y\in\left\langle{\gamma^{s-1}}\right\rangle\setminus\left\{{0}\right\}}\xi^{\operatorname{Tr}(y)}=-1&\text{if }J=S,\\ \sum_{y\in\mathcal{R}}\xi^{\operatorname{Tr}(y)}-\sum_{y\in\left\langle{\gamma}\right\rangle}\xi^{\operatorname{Tr}(y)}=0&\text{if }J=U,\\ \sum_{y\in\left\langle{\gamma}\right\rangle}\xi^{\operatorname{Tr}(y)}-\sum_{y\in\left\langle{\gamma^{s-1}}\right\rangle}\xi^{\operatorname{Tr}(y)}=0&\text{if }J=R.\end{cases}$
Instead, if $J=U$, for $I\in\left\{{Z,U}\right\}$ we already discussed the value of (86) above. In this case, it remains to consider $I=S$ and $I=R$. In the first case, for any $a_{k}\in I$, we have $a_{k}J=I$. However, as the sum runs over $J$, we observe every element in $a_{k}J=I$ exactly $\left|\,{U}\,\right|/\left|\,{S}\,\right|=q^{s-1}$ many times and hence,
(91) $\sum_{x_{k}\in U}\xi^{\operatorname{Tr}(a_{k}x_{k})}=q^{s-1}\sum_{y\in S}\xi^{\operatorname{Tr}(y)}=q^{s-1}\cdot(-1)=-q^{s-1}.$
In the latter case (where $I=R$), consider any $a_{k}\in I$. Then, $a_{k}J=\left\langle{\gamma^{i}}\right\rangle\setminus\left\langle{\gamma^{i+1}}\right\rangle$. As $\sum_{y\in\left\langle{\gamma^{i}}\right\rangle\setminus\left\langle{\gamma^{i+1}}\right\rangle}\xi^{\operatorname{Tr}(y)}=0$, we also observe
(92) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=0.$
Lastly, let us consider the case where $I=J=R$. If $a_{k}\in\langle\gamma^{i}\rangle\setminus\langle\gamma^{i+1}\rangle$, then $a_{k}R=(\langle\gamma^{i+1}\rangle\setminus\{0\})^{w}\times(\langle\gamma^{i+1}\rangle)^{v}$, depending on whether the product $a_{k}x_{k}$ can be zero. Note that there are $|\langle\gamma^{s-i}\rangle\setminus\langle\gamma^{s-1}\rangle|=q^{i}-q$ many $x_{k}\in R$ which make the product zero. Thus, the ideal $\langle\gamma^{i+1}\rangle$ of size $q^{s-i-1}$ is seen $v=q^{i}-q$ many times and we are left with $|R|-(q^{i}-q)q^{s-i-1}=q^{s-i}-q$ many $x_{k}\in R$, which see $\langle\gamma^{i+1}\rangle\setminus\{0\}$ exactly $w=(q^{s-i}-q)/(q^{s-i-1}-1)=q$ times. Thus, we get that
(93) $\sum_{x_{k}\in R}\xi^{\operatorname{Tr}{(x_{k}a_{k})}}=(q^{i}-q)\sum_{y\in\langle\gamma^{i+1}\rangle}\xi^{\operatorname{Tr}{(y)}}+q\sum_{y\in\langle\gamma^{i+1}\rangle\setminus\{0\}}\xi^{\operatorname{Tr}{(y)}}=(q^{i}-q)\cdot 0+q\cdot(-1)=-q.$
Table 1 gives an overview of the values of the sum in (86) for all $I,J\in\left\{{Z,U,S,R}\right\}$. It is worth noting that these values are independent of the choice of $a_{k}\in I$. This independence yields to a well-defined expression for the Krawtchouk coefficients, wherefore they exist. By repeating the procedure for each position $k$ and summing over all possible compositions $t\in\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho)$, we obtain the desired result.
Table 1. Values of the expression $\sum_{x_{\ell}\in J}\xi^{\operatorname{Tr}(a_{k}\cdot x_{\ell})}$ for any $I,J\in\left\{{Z,U,S,R}\right\}$ as given in (86).
$I$ $J$ $Z$ $U$ $S$ $R$
$Z$ 1 $q^{s-1}(q-1)$ $q-1$ $q^{s-1}-q$
$U$ 1 0 $-1$ 0
$S$ 1 $-q^{s-1}$ $q-1$ $q^{s-1}-q$
$R$ 1 0 $q-1$ $-q$
Let us consider an example to get an idea of the formula stated above.
Example 31.
Consider again the code $\mathcal{C}\subset(\mathbb{Z}/{9}\mathbb{Z})^{3}$ generated by $G=\begin{pmatrix}3&2&8\end{pmatrix}$. Notice, that there are three distinct homogeneous weight-unit decompositions occurring in $\mathcal{C}$, namely,
(94) $\pi^{(0)}=(3,0,0,0),\quad\pi^{(1)}=(1,0,2,0),\quad\pi^{(2)}=(0,2,1,0),$
with the corresponding homogeneous weight-unit enumerators given by
(95) $\mathcal{D}_{\pi^{(0)}}^{\mathsf{Hom}}(\mathcal{C})=1,\quad\mathcal{D}_{\pi^{(1)}}^{\mathsf{Hom}}(\mathcal{C})=2,\quad\mathcal{D}_{\pi^{(2)}}^{\mathsf{Hom}}(\mathcal{C})=6.$
Note that the dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ is generated by
(96) $G^{\perp}=\begin{pmatrix}1&0&3\\ 0&1&2\end{pmatrix}.$
Let $\rho=(2,0,1,0)\in\mathbb{D}_{9,3}^{\mathsf{Hom}}$ be a homogeneous weight-unit decomposition in $(\mathbb{Z}/{9}\mathbb{Z})^{3}$. By calculating the codewords of $\mathcal{C}^{\perp}$, we observe that $\mathcal{D}_{\rho}^{\mathsf{Hom}}(\mathcal{C}^{\perp})=2$. For every $\pi\in\left\{{\pi^{(0)},\pi^{(1)},\pi^{(2)}}\right\}$ let us now compute the Krawtchouk coefficient $K^{\mathsf{Hom}}_{\pi}(\rho)$ as stated in Theorem 30.
Let us start by choosing $\pi=\pi^{(0)}$. Notice that there is only one composition $\rho$ with respect to $\pi^{(0)}$, i.e.,
(97) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi^{(0)}}(\rho)=\Big{\{}(\underbrace{2,0,1,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{0,0,0,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}})\Big{\}}.$
Hence, we obtain
(98) $K^{\mathsf{Hom}}_{\pi^{(0)}}(\rho)=\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}=\binom{\pi_{Z}}{t_{Z}}(p-1)^{t_{ZS}}=\binom{3}{2,0,1,0}2^{1}=6.$
Now let $\pi=\pi^{(1)}$. Note that the set of compositions $\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho)$ of $\rho$ with respect to $\pi$ consists of only two elements, i.e.,
(99) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi^{(1)}}(\rho)=\Big{\{}t=(\underbrace{1,0,0,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{1,0,1,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}}),s=(\underbrace{0,0,1,0}_{s_{Z}},\underbrace{0,0,0,0}_{s_{U}},\underbrace{2,0,0,0}_{s_{S}},\underbrace{0,0,0,0}_{s_{R}})\Big{\}}.$
With this we obtain
(100) $K^{\mathsf{Hom}}_{\pi^{(1)}}(\rho)$ $=\left(\binom{\pi_{Z}}{t_{Z}}\binom{\pi_{S}}{t_{S}}(p-1)^{t_{SS}}\right)+\left(\binom{\pi_{Z}}{s_{Z}}\binom{\pi_{S}}{s_{S}}(p-1)^{s_{SS}}\right)$
(101) $=\binom{1}{1,0,0,0}\binom{2}{1,0,1,0}2^{1}+\binom{1}{0,0,1,0}\binom{2}{2,0,0,0}2^{1}$
(102) $=6.$
Lastly, for $\pi^{(2)}$ we have
(103) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi^{(2)}}(\rho)=\Big{\{}t=(\underbrace{0,0,0,0}_{t_{Z}},\underbrace{2,0,0,0}_{t_{U}},\underbrace{0,0,1,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}}),\ s=(\underbrace{0,0,0,0}_{s_{Z}},\underbrace{1,0,1,0}_{s_{U}},\underbrace{1,0,0,0}_{s_{S}},\underbrace{0,0,0,0}_{s_{R}})\Big{\}},$
which yields to
(104) $K^{\mathsf{Hom}}_{\pi^{(2)}}(\rho)$ $=\left(\binom{\pi_{U}}{t_{U}}\binom{\pi_{S}}{t_{S}}(p-1)^{t_{SS}}\right)+\left(\binom{\pi_{U}}{s_{U}}\binom{\pi_{S}}{s_{S}}(-1)^{s_{US}}\right)$
(105) $=\binom{2}{2,0,0,0}\binom{1}{0,0,1,0}2^{1}+\binom{2}{1,0,1,0}\binom{1}{1,0,0,0}(-1)^{1}$
(106) $=0.$
Finally, we apply Equation (82) and we indeed obtain
(107) $\mathcal{D}_{\rho}^{\mathsf{Hom}}(\mathcal{C}^{\perp})$ $=\frac{1}{9}\left(K^{\mathsf{Hom}}_{\pi^{(0)}}(\rho)\mathcal{D}_{\pi^{(0)}}^{\mathsf{Hom}}+K^{\mathsf{Hom}}_{\pi^{(1)}}(\rho)\mathcal{D}_{\pi^{(1)}}^{\mathsf{Hom}}+K^{\mathsf{Hom}}_{\pi^{(2)}}(\rho)\mathcal{D}_{\pi^{(2)}}^{\mathsf{Hom}}\right)$
(108) $=\frac{1}{9}(6\cdot 1+6\cdot 2+0\cdot 6)=2.$
Finally, note that the homogeneous unit-weight partition (see Definition 29) defines the coarsest partition that allows us to obtain MacWilliams-type identities for the homogeneous metric. In fact, we are looking for a partition of the code that gives well-defined Krawtchouk coefficients and is finer than the homogeneous-weight partition, as we want to recover the weight enumerator.
3.3. The Subfield Metric
Let us focus on the subfield metric defined over the finite field $\mathbb{F}_{p^{r}}$. Since $\mathbb{F}_{p^{r}}$ is a finite chain ring with $s=1$ and the subfield weight is additive, using the Lee partition allows us to obtain MacWilliams-type identities for the subfield weight. However, this is not the coarsest partition we can choose. Again, a natural choice is to consider the subfield decomposition of the codewords defined as follows: given a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ the subfield decomposition of a codeword $c\in\mathcal{C}$ is a triple $(s_{0}(c),s_{1}(c),s_{\lambda}(c))$, where for each $i\in\left\{{0,1,\lambda}\right\}$ we define
(109) $s_{i}(c):=\left|\,{\left\{{j=1,\ldots,n\;|\;\operatorname{wt}_{\lambda}(c_{j})=i}\right\}}\,\right|.$
We then can partition the code into blocks containing the codewords with the same subfield weight decomposition. The sizes of the blocks determine the subfield decomposition enumerator of a code $\mathcal{C}$ and we denote it by $\lambda\mathsf{dec}_{\mathcal{C}}$.
Example 32.
Consider the finite field $\mathbb{F}_{8}=\mathbb{F}_{2}[\alpha]$ where $\alpha^{3}=\alpha+1$ and the two codes $\mathcal{C}_{1}=\left\langle{\begin{pmatrix}1,\alpha,\alpha^{2}\end{pmatrix}}\right\rangle$ and $\mathcal{C}_{2}=\left\langle{\begin{pmatrix}\alpha^{2}+2,2\alpha+1,\alpha^{2}+1\end{pmatrix}}\right\rangle$. We denote by $\mathcal{C}^{\perp}_{1}$ and $\mathcal{C}^{\perp}_{2}$ the dual of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Table 2 then shows that $\lambda\mathsf{dec}_{\mathcal{C}_{1}}=\lambda\mathsf{dec}_{\mathcal{C}_{2}}$, but $\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{1}}\neq\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{2}}$.
Table 2. Subfield weight enumerator for the codes $\mathcal{C}_{1},\mathcal{C}_{2}$ and their respective duals $\mathcal{C}^{\perp}_{1}$ and $\mathcal{C}^{\perp}_{2}$.
subfield weight decomp. $\lambda\mathsf{dec}_{\mathcal{C}_{1}}$ $\lambda\mathsf{dec}_{\mathcal{C}_{2}}$ $\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{1}}$ $\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{2}}$
(3, 0, 0) 1 1 1 1
(0, 3, 0) 0 0 0 0
(0, 0, 3) 4 4 26 27
(2, 1, 0) 0 0 0 0
(2, 0, 1) 0 0 0 0
(1, 2, 0) 0 0 0 0
(1, 0, 2) 0 0 15 15
(0, 2, 1) 0 0 0 3
(0, 1, 2) 3 3 15 12
(1, 1, 1) 0 0 6 6
Note that Example 32 also shows that partitioning a code $\mathcal{C}$ into codewords of the same subfield weight does not yield to MacWilliams identities. Indeed, for $\lambda>1$, Example 32 shows that the subfield weight enumerators of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ coincide (i.e., $\lambda\mathsf{we}_{\mathcal{C}_{1}}=\lambda\mathsf{we}_{\mathcal{C}_{2}}$), whereas the subfield weight enumerators of their duals are different (i.e, $\lambda\mathsf{we}_{\mathcal{C}^{\perp}_{1}}\neq\lambda\mathsf{we}_{\mathcal{C}^{\perp}_{2}}$).
Remark 33.
For the MacWilliams-type identities for the subfield weight we consider a character of the form $\chi_{a}(x)=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}$ of $\mathbb{F}_{p^{r}}^{n}$. For quantum error-correcting codes, the use of the alternating or Hermitian form instead of the standard inner product are usually a more natural choice (see [8, 22] for more details).
However, the Hermitian form for any two vectors $x,y\in\mathbb{F}_{p^{r}}^{n}$ is defined for $r=2$, for which the MacWilliams identities have been shown to exist [22]. The alternating inner product of two vectors $x,y\in\mathbb{F}_{p^{r}}^{n}$ is defined only for an even $r$. We fix a $\mathbb{F}_{p}$-basis $\Gamma=\{1,\gamma,\ldots,\gamma^{r-1}\}$ of $\mathbb{F}_{p^{r}}$ and expand the vectors over the base field. That is,for any coordinate $1\leq i\leq n$
(110) $x_{i}=x_{i}^{(0)}+x_{i}^{(1)}\gamma+\ldots+x_{i}^{(r-1)}\gamma^{r-1}\in\mathbb{F}_{p^{r}}$
and we write $\Gamma(x_{i})=(x_{i}^{(0)},\ldots,x_{i}^{(s-1)})\in\mathbb{F}_{p^{r}}.$ Then, we denote by $x\star y=\sum_{i=1}^{n}x_{i}\star y_{i}$ the alternating inner product of $x,y\in\mathbb{F}_{p^{r}}$ where
(111) $x_{i}\star y_{i}=\sum_{j=0}^{r/2-1}x_{i}^{(j)}y_{i}^{(j+r/2)}-x_{i}^{(j+r/2)}y_{i}^{(j)}.$
Then a character of $\mathbb{F}_{p^{r}}$ has the form $\chi_{a}(x)=\xi^{\operatorname{Tr}(a\star x)}$. Notice that finding a partition $\mathcal{P}=P_{1}\;|\;\ldots\;|\;P_{t}$ of $\mathbb{F}_{p^{r}}^{n}$ that allows for an expression of $\sum_{x\in P_{j}}\xi^{\operatorname{Tr}(a\star x)}$ that is independent on the choice of $a\in P_{i}$ is way more involved than for the case where we consider the standard inner product. In fact, to get a well-defined Krawtchouk coefficient one would need to go through all possible $\Gamma(x),$ leaving us with $p^{rn}$ choices.
Another idea, similar to the Lee partition, is to consider an orbit partition for which, by [19], MacWilliams identities hold. Based on the definition of the subfield weight over $\mathbb{F}_{p^{r}}$, we consider the multiplicative group action of $\mathbb{F}_{p}^{\times}$ on $\mathbb{F}_{p^{r}}$. We partition the finite field $\mathbb{F}_{p^{r}}$ into distinct orbits of multiplication with $\mathbb{F}_{p}^{\times}$, i.e., $I_{\alpha}:=\alpha\mathbb{F}_{p}^{\times}$, where $\alpha\in\mathbb{F}_{p^{r}}$. To keep elements of different subfield weight separate, it is natural to consider the orbits $I_{0}:=\left\{{0}\right\}$ and $I_{1}:=\mathbb{F}_{p}^{\times}$. In the following we denote $\alpha_{0}:=0$ and $\alpha_{1}:=1$ and write $I_{0}=\alpha_{0}\mathbb{F}_{p}^{\times}$ and $I_{1}=\alpha_{1}\mathbb{F}_{p}^{\times}$. Then, for $i=2,\ldots,\frac{p^{r}-1}{p-1}$, we recursively choose an element $\alpha_{i}\in\mathbb{F}_{p^{r}}\setminus\left(\bigcup_{j=0}^{i-1}I_{j}\right)$ and define $I_{i}:=\alpha_{i}\mathbb{F}_{p}^{\times}.$ Defining $\ell:=\frac{p^{r}-1}{p-1}$, we get that
(112) $\mathcal{I}=I_{0}\;|\;\ldots\;|\;I_{\ell},$
forms a partition of $\mathbb{F}_{p^{r}}$ which we call the subfield-orbit partition of $\mathbb{F}_{p^{r}}$. The subfield-orbit partition $\mathcal{I}^{\lambda}$ allows to define the subfield-orbit decomposition of a tuple $x\in\mathbb{F}_{p^{r}}^{n}$. That is, given a vector $x\in\mathbb{F}_{p^{r}}^{n}$ we define its subfield-orbit decomposition to be the $(\ell+1)$-tuple $\sigma(x)=(\sigma_{0}(x),\sigma_{1}(x),\ldots,\sigma_{\ell}(x))$ where, for every $i\in\left\{{0,\ldots,\ell}\right\}$,
(113) $\sigma_{i}(x):=\left|\,{\left\{{k=1,\ldots,n\;|\;x_{k}\in I_{i}}\right\}}\,\right|.$
Given a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$, we can partition the codewords of $\mathcal{C}$ via their subfield-orbit partition. This is a partition of $\mathcal{C}$ given by the orbits of $\mathbb{F}_{p}^{\times}$ which form a subgroup of the automorphism group of $\mathbb{F}_{p^{r}}$ and, hence, we obtain MacWilliams identities for the subfield-orbit enumerator of the code $\mathcal{C}$. Additionally, note that every $c\in P^{\lambda}_{\pi}$ has the same subfield weight given by $\operatorname{wt}_{\lambda}(c)=\sigma_{1}+\lambda\sum_{i=2}^{\ell}\sigma_{i}$. Hence, the subfield weight enumerator $\lambda\mathsf{we}_{\mathcal{C}}$ can immediately be derived from the subfield-orbit enumerator.
The subfield-orbit partition of $\mathbb{F}_{p^{r}}^{n}$, however, is not the coarsest partition that recovers the subfield weight enumerator and allows for MacWilliams-type identities. In fact, we can do better by defining a reflexive partition consisting of only four blocks that is not based on the orbits of a subgroup. More explicitly, with an eye on Example 11, we will distinguish the elements in $\mathbb{F}_{p^{r}}^{n}$ not only by their subfield weight, but also by their trace. As we still want to be able to recover the subfield weight of a vector, we subpartition the elements of subfield weight $\lambda$ (i.e., the elements in $\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}$) into two sets: the first one contains all such elements with trace equal to 00 and the second one contains all remaining elements. That is, we use $\mathcal{I}^{\lambda}=I^{\lambda}_{0}\;|\;I^{\lambda}_{1}\;|\;I^{\lambda}_{2}\;|\;I^{\lambda}_{3}$ to partition $\mathbb{F}_{p^{r}}$ where
(114) $I^{\lambda}_{0}:=\left\{{0}\right\},\quad I^{\lambda}_{1}:=\mathbb{F}_{p}^{\times},\quad I^{\lambda}_{2}:=\left\{{x\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}\;|\;\operatorname{Tr}(x)=0}\right\}\quad\text{and}\quad I^{\lambda}_{3}:=\left\{{x\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}\;|\;\operatorname{Tr}(x)\neq 0}\right\}.$
We call $\mathcal{I}^{\lambda}$ the subfield-trace partition of $\mathbb{F}_{p^{r}}$. Recall that the trace function is uniformly distributed among $\mathbb{F}_{p}$ and that for each $x\in\mathbb{F}_{p}$ we have $\operatorname{Tr}(x)=mx\mod p$. Hence, the following is a simple consequence.
Corollary 34.
[…]
To extend the partition to $\mathbb{F}_{p^{r}}^{n}$ we first define the subfield-trace decomposition of an element $x\in\mathbb{F}_{p^{r}}^{n}$.
Definition 35.
Given a finite field $\mathbb{F}_{p^{r}}^{n}$ and an element $x\in\mathbb{F}_{p^{r}}^{n}$. The subfield-trace decomposition of $x$ is a $4$-tuple $\pi^{\mathsf{\lambda}}(x)=(\pi^{\mathsf{\lambda}}_{0}(x),\pi^{\mathsf{\lambda}}_{1}(x),\pi^{\mathsf{\lambda}}_{2}(x),\pi^{\mathsf{\lambda}}_{3}(x))$ where, for each $i\in\left\{{0,\ldots,3}\right\}$, we have
(115) $\pi^{\mathsf{\lambda}}_{i}(x)=\left|\,{\left\{{j=1,\ldots,n\;|\;x_{j}\in I^{\lambda}_{i}}\right\}}\,\right|.$
Similar to the Lee and homogeneous metric, we denote the set of all subfield-trace decompositions in $\mathbb{F}_{p^{r}}^{n}$ by
(116) $\mathbb{D}^{\lambda}_{p,r,1,n}=\left\{{\pi\in\left\{{0,\ldots,n}\right\}^{4}\;\big{|}\;\sum_{i=0}^{3}\pi_{i}=n}\right\}.$
Given that $\mathbb{D}^{\lambda}_{p,r,1,n}=\left\{{\pi^{(0)},\ldots,\pi^{(D-1)}}\right\}$, this induces a subfield-trace partition $\mathcal{P}^{\lambda}$ of $\mathbb{F}_{p^{r}}^{n}$ into the vectors of the same subfield-trace decomposition, i.e.,
(117) $\mathcal{P}^{\lambda}:=P^{\lambda}_{\pi^{(0)}}\;|\;\ldots\;|\;P^{\lambda}_{\pi^{(D-1)}}.$
Hence, for a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$, the number of codewords $c\in\mathcal{C}$ of subfield-trace decomposition $\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}$ is given by
(118) $\mathcal{D}_{\pi}^{\lambda}(\mathcal{C}):=\left|\,{P^{\lambda}_{\pi}\cap\mathcal{C}}\,\right|.$
Note also here that every codeword $c\in\mathcal{C}$ with $c\in P^{\lambda}_{\pi}$ has the same subfield weight which is calculated by $\operatorname{wt}_{\lambda}(c)=\pi_{1}+(\pi_{2}+\pi_{3})\lambda$. Thus, the subfield-weight enumerator $\lambda\mathsf{dec}_{\mathcal{C}}$ of a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ can immediately be derived from the subfield-trace enumerator of $\mathcal{C}$.
We are now able to state the MacWilliams identities for the subfield-trace enumerator.
Theorem 36 (MacWilliams-type Identities for the Subfield-Trace Enumerator).
Consider a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given $\rho\in\mathbb{D}^{\lambda}_{p,r,1,n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of subfield-trace decomposition $\rho$ is given by
(119) $\mathcal{D}_{\rho}^{\lambda}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C}).$
Here $K^{\lambda}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$, which for a given $p$-th root of unity $\xi$ is defined as
(120) $K^{\lambda}_{\pi}(\rho)=\sum_{t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)}\left(\prod_{i=0}^{3}\binom{\pi_{i}}{t_{i0},\ldots,t_{i3}}\right)\cdot\kappa,$
where
(121) $\kappa=\begin{cases}(p-1)^{t_{01}+t_{11}+t_{21}+t_{03}}(-1)^{t_{31}}(-p)^{t_{22}}(p^{r-1}-p)^{t_{02}+t_{12}}(-p^{r-1})^{t_{13}}(p^{r-1})^{t_{03}}\mathbbm{1}_{t_{23}=t_{32}=t_{33}=0}&\text{if }r\;|\;p,\\ (-1)^{t_{11}+t_{22}+t_{33}+t_{32}}(p-1)^{t_{01}+t_{03}}(1-p)^{t_{2}3}(p^{r-1}-1)^{t_{02}+t_{03}+t_{12}+t_{21}}(1-p^{r-1})^{t_{13}}&\text{otherwise}.\end{cases}$
Proof.
The multinomial coefficient appearing in the expression for the Krawtchouk coefficient is derived as in the proof of Theorem 30. Thus, it suffices to show that, for any $i\in\left\{{0,\ldots,3}\right\}$, $j\in\left\{{0,\ldots,3}\right\}$ and $a,b\in I^{\lambda}_{j}$ with $a\neq b$ it holds that
(122) $\sum_{x\in I^{\lambda}_{i}}\xi^{\operatorname{Tr}(ax)}=\sum_{x\in I^{\lambda}_{i}}\xi^{\operatorname{Tr}(bx)}.$
That is, in accordance with Example 11, number 3., we want to show
(123) $\sum_{x\in I^{\lambda}_{i}}\chi_{a}(x)=\sum_{x\in I^{\lambda}_{i}}\chi_{b}(x).$
Therefore, we compute $\sum_{x\in I^{\lambda}_{i}}\chi_{a}(x)$ in each case and we show that the sum is independent on the choice of $a\in I^{\lambda}_{i}$. For $a=0$ and for every $i\in\left\{{0,\ldots,3}\right\}$ we observe that
(124) $\sum_{x\in I^{\lambda}_{i}}\chi_{a}(x)=|I^{\lambda}_{i}|,$
where the values of $\left|\,{I^{\lambda}_{i}}\,\right|$, for $i=2,3$, can be found in Corollary 34. Similarly, for any $j\in\left\{{0,\ldots,3}\right\}$ and $a\in I^{\lambda}_{j}$ we have
(125) $\sum_{x\in I^{\lambda}_{0}}\chi_{a}(x)=1.$
For the remaining cases, and we make use of Lemma 13 and 14 and we identify whether the character $\chi_{a}(x)$ is generating for the subgroup $\mathbb{F}_{p}\subset\mathbb{F}_{p^{r}}$. Recall that, a character $\chi_{a}$ of $\mathbb{F}_{p^{r}}$ is generating for the subgroup $H$ if its kernel is contained in $H$.
Consider first $a\in I^{\lambda}_{1}$, that is $a\in\mathbb{F}_{p}^{\times}$. By $\mathbb{F}_{p}$-linearity of the trace function, we have $\operatorname{Tr}(ax)=a\operatorname{Tr}(x)$ for any $x\in\mathbb{F}_{p^{r}}$. If $r\;|\;p$, for any $x\in\mathbb{F}_{p}$ we have $\operatorname{Tr}(x)=0$, i.e., $\chi_{a}$ is generating for $\mathbb{F}_{p}$. Hence, by Lemma 13, $\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=p$. Therefore,
(126) $\sum_{x\in\mathbb{F}_{p}^{\times}}\chi_{a}(x)=p-1.$
[…]
(127) $\sum_{x\in\mathbb{F}_{p}^{\times}}\chi_{a}(x)=-1.$
[…]
(128) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=|I^{\lambda}_{2}|=p^{r-1}-p&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=|I^{\lambda}_{2}|=p^{r-1}-1&\text{otherwise}.\end{cases}$
Let $a\in I^{\lambda}_{2}$, i.e., we have $\operatorname{Tr}(a)=0$. From the $\mathbb{F}_{p}$-linearity of the trace we obtain that $\operatorname{Tr}(ax)=\operatorname{Tr}(a)x=0$ for any $a\in I^{\lambda}_{2}$ and hence, $\chi_{a}$ is generating for $\mathbb{F}_{p}$ for any $a$ Then, by Lemma 13, we get that
(129) $\sum_{x\in\mathbb{F}_{p}^{\times}}\chi_{a}(x)=\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)-\chi_{a}(0)=p-1.$
[…]
(130) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=0-p=-p&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=0-1=-1&\text{otherwise}.\end{cases}$
Finally, we assume that $a\in I^{\lambda}_{3}$. Note that, if $x=1\in\mathbb{F}_{p}$, then $\operatorname{Tr}(ax)=\operatorname{Tr}(a)\neq 0$. Hence, $\chi_{a}$ is non-generating for $\mathbb{F}_{p}$ over $\mathbb{F}_{p^{r}}$ which implies (using Corollary 14) that
(131) $\sum_{x\in\mathbb{F}_{p}^{\times}}\chi_{a}(x)=-1.$
Similarly to the previous cases, we obtain
(132) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=0-0=0&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=0-1=-1&\text{otherwise}.\end{cases}$
Observing that $\sum_{x\in I^{\lambda}_{3}}\chi_{a}(x)=\sum_{x\in\mathbb{F}_{p^{r}}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)$, yields the last cases for each $a$.
[…]
00 $1$ $2$ $3$
00 $1$ $1$ $1$ $1$
$1$ $p-1$ $p-1$ $p-1$ $-1$
$2$ $p^{r-1}-p$ $p^{r-1}-p$ $-p$ 00
$3$ $p^{r-1}(p-1)$ $-p^{r-1}$ 00 00
(a) Values of $\sum_{x\in I^{\lambda}_{j}}\chi_{a}$ for $a\in I^{\lambda}_{i}$, $i,j\in\left\{{0,\ldots,3}\right\}$ and $r\;|\;p$.
00 $1$ $2$ $3$
00 $1$ $1$ $1$ $1$
$1$ $p-1$ $-1$ $p-1$ $-1$
$2$ $p^{r-1}-1$ $p^{r-1}-1$ $-1$ $-1$
$3$ $(p^{r-1}-1)(p-1)$ $-(p^{r-1}-1)$ $1-p$ $1$
[…]
Let us give an example to illustrate the calculations.
Example 37.
Given the finite field $\mathbb{F}_{27}=\mathbb{F}_{3}[\alpha]$ where $\alpha^{3}=\alpha+2$. Using the subfield-trace partition (as stated in (114)) we partition $\mathbb{F}_{27}$ into the following four blocks
(133) $I^{\lambda}_{0}$ =$=\left\{{0}\right\},\quad I^{\lambda}_{1}=\mathbb{F}_{27}^{\times}=\left\{{1,2}\right\},\quad I^{\lambda}_{2}=\left\{{\alpha,\alpha+1,\alpha+2,2\alpha,2\alpha+1,2\alpha+2}\right\},\quad\text{and}$
(134) $I^{\lambda}_{3}$ =$=\{\alpha^{2},\alpha^{2}+1,\alpha^{2}+2,\alpha^{2}+\alpha,\alpha^{2}+\alpha+1,\alpha^{2}+\alpha+2,\alpha^{2}+2\alpha,\alpha^{2}+2\alpha+1,\alpha^{2}+2\alpha+2,$
(135) $\qquad 2\alpha^{2},2\alpha^{2}+1,2\alpha^{2}+2,2\alpha^{2}+\alpha,2\alpha^{2}+\alpha+1,2\alpha^{2}+\alpha+2,2\alpha^{2}+2\alpha,2\alpha^{2}+2\alpha+1,2\alpha^{2}+2\alpha+2\}.$
Consider the code $\mathcal{C}\subset\mathbb{F}_{27}^{3}$ generated by $G=\begin{pmatrix}1&\alpha&1\end{pmatrix}$. By computing the codewords of $\mathcal{C}$ we observe that they can be partitioned into six distinct subfield-trace decompositions, namely $(3,0,0,0),(0,2,1,0),(0,0,2,1),(0,0,1,2),(0,1,0,2)$ and $(0,0,0,3)$. Each of the six subfield-trace decompositions are represented by $1,2,6,4,2$ and $12$ codewords, respectively.
The dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ has a generator matrix of the form
(136) $G^{\perp}=\begin{pmatrix}1&0&2\\ 0&1&\alpha\end{pmatrix}.$
It is easily checked that the vectors $(1,0,2)$ and $(2,0,1)$ are the only two codewords of $\mathcal{C}^{\perp}$ of subfield-trace decomposition equal to $\rho=(1,2,0,0)$. Thus, $\mathcal{D}_{\rho}^{\lambda}(\mathcal{C}^{\perp})=2$.
We now verify that indeed $\frac{1}{|\mathcal{C}|}\sum_{\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}(\rho)}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=2$. We show detailed calculations of the Krawtchouk coefficient only for $\pi^{(0)}=(3,0,0,0)$ and $\pi^{(1)}=(0,2,1,0)$. Let us first assume, that $\pi=\pi^{(0)}$. Similar to the examples of the Lee and the homogeneous metric, we observe that
(137) $\mathrm{Comp}^{\lambda}_{\pi}(\rho)=\{(\underbrace{1,2,0,0}_{\pi_{0}},\underbrace{0,\ldots,0}_{\pi_{1}},\underbrace{0,\ldots,0}_{\pi_{2}},\underbrace{0,\ldots,0}_{\pi_{3}})\}.$
That is, for $t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$, we have $t_{00}=1,t_{01}=2$ and $t_{02}=\ldots=t_{33}=0$. This yields
(138) $K^{\lambda}_{\pi}(\rho)=\binom{3}{2,1}(3-1)^{2}=12,\quad\text{and}\quad K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=12\cdot 1=12.$
Consider now $\pi=\pi^{(1)}$. For visual reasons, Figure 2 shows the set $\mathrm{Comp}^{\lambda}_{\pi}(\rho)$ of compositions of $\rho$ with respect to $\pi$.
[…]
Figure 2. Illustration of $\mathrm{Comp}^{\lambda}_{\pi}(\rho)$.
For the first option $t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$ shown in Figure 2, we have $t_{01}=t_{11}=t_{21}=1$ and all other $t_{ij}$’s are equal to 00. The second option, i.e., $s\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$, consists of $s_{11}=2$ and $s_{20}=1$ and $s_{ij}=0$ for all other indices. Hence,
(139) $K^{\lambda}_{\pi}(\rho)=\binom{2}{1,1}\binom{1}{1}(3-1)^{1+1}+\binom{2}{2}\binom{1}{1}(3-1)^{2}=12,\quad\text{and}\quad K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=24.$
The calculations of the remaining Krawtchouk coefficients are similar to the calculations of $K^{\lambda}_{\pi^{(0)}}(\rho)$ and $K^{\lambda}_{\pi^{(1)}}(\rho)$ leading to
(140) $\frac{1}{|\mathcal{C}|}\sum_{\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}(\rho)}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=\frac{1}{27}\left(12+24+0\cdot 6+(-3)\cdot 4+(-3)\cdot 2+3\cdot 12\right)=\frac{1}{27}\cdot 54=2.$
Finally, we want to highlight at this point that the subfield-trace partition (see (117)) is indeed the coarsest partition that allows for the MacWilliams identities to exist and to recover the subfield weight enumerator.
4. Linear Programming Bound
One of the tightest bounds for the size of linear codes is given by the Linear Programming (LP) bound. Such bounds have been introduced by Delsarte [16] and usually stem from association schemes. However, given the MacWilliams identities, one can easily derive an LP bound. For the homogeneous and the subfield metric no such LP bound has been stated. An exception is the Lee metric, where an association scheme using the Lee decomposition has been employed to derive a Lee-metric LP bound [6, 41]. For the homogeneous metric over a Frobenius ring a different approach to derive an LP bound has been presented in [11]. In this Section, we derive LP bounds stemming from the MacWilliams identities presented in Section 3.
Given a finite chain ring $\mathcal{R}$, an additive weight $\mathrm{wt}:\mathcal{R}\to\mathbb{Q}$ and an $\mathcal{R}$-linear code $\mathcal{C}$, recall the definition of the decomposition enumerator of $\mathcal{C}$
(141) $\mathcal{D}_{\pi}^{\operatorname{wt}}(\mathcal{C})=\left|\,{\left\{{c\in\mathcal{C}\;|\;\pi^{\operatorname{wt}}(c)=\pi}\right\}}\,\right|$
with respect to the weight $\operatorname{wt}$ (see (51), (81) and (118) for the Lee, homogeneous-unit and subfield-trace enumerator, respectively). Moreover, if $\left|\,{\mathbb{D}^{\operatorname{wt}}_{p,r,s,n}}\,\right|=D$, we denote by $\pi^{(0)}=(n,0,\dots,0)$ the decomposition of the all-zero vector and by $\pi^{(1)},\dots,\pi^{(D-1)}$ the other decompositions in $\mathbb{D}^{\operatorname{wt}}_{p,r,s,n}$.
The MacWilliams-type identities for the Lee decomposition enumerator, given in Corollary 26, allow us to straightforwardly derive an LP bound for any $\mathcal{R}$-linear code and any additive weight. Moreover, the MacWilliams-type identities given in Theorem 30 and Theorem 36 allow to derive an LP bound for the homogeneous and subfield metric respectively, which is less complex (i.e., that involves fewer variables).
From now on, $B_{p,r,s,n}(n,d_{\operatorname{wt}})$ denotes the maximum number of codewords of an $\mathcal{R}$-linear code of length $n$ and minimum $\mathrm{wt}$-distance $d_{\operatorname{wt}}$.
Theorem 38 (Linear Programming Bound).
For a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ of minimum distance $d_{\operatorname{wt}}$ it holds that
(142) $B_{p,r,s,n}(n,d_{\operatorname{wt}})\leq\max\left(\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})\right),$
where the maximum is taken over all $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})$ subject to the following conditions:
  • $\mathcal{D}_{\pi^{(0)}}^{\operatorname{wt}}(\mathcal{C})=1$;
  • $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})\geq 0$ for all $11\leq i\leq D-$ 1 and $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})=0$ for all $i\in I$;
  • $\sum_{i=1}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})K_{\pi^{(j)}}(\pi^{(i)})\geq 0$,
where $I=\{i\ \;|\;\ x\in\mathcal{R}^{n}\text{ such that }\pi(x)=\pi^{(i)}\text{ and }\mathrm{wt}(x)<d_{\operatorname{wt}}\}$.
If $\operatorname{wt}$ is an additive weight over $\mathcal{R}$ and $\mathcal{D}_{\pi}^{\operatorname{wt}}(\mathcal{C})$ is the Lee partition enumerator, the LP bound requires a large number of variables. Thus, the program cannot run efficiently and is not computationally feasible.
4.1. Lee metric
The Lee scheme, introduced in [6] and investigated in [4, 5, 41], provides a framework for studying the maximum size of a linear Lee-metric code. In the Lee-metric case the LP bound Theorem 38 can be refined as follows.
Through this section, let $\mathcal{R}=(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and let $M$ be the maximum value of the Lee-weight in $\mathbb{Z}/{p^{s}}\mathbb{Z}$. Given a $\mathbb{Z}/{p^{s}}\mathbb{Z}$-linear code $\mathcal{C}$ of length $n$, let $c\in\mathcal{C}$ denote a codeword. In the Hamming metric, whenever we multiply each component of $c$ by a unit $r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$, we obtain a codeword $r\cdot c$ of the same Hamming weight as $c$. This property does not hold in the Lee metric. However, the Lee decompositions of $c$ and $r\cdot c$ are related. Indeed, the Lee decomposition of $r\cdot c$ is a permutation of the Lee decomposition of $c$. That is, if $\pi^{\mathsf{L}}(c)=(\pi^{\mathsf{L}}_{0},\dots,\pi^{\mathsf{L}}_{M})$ is the Lee decomposition of $c$, then the Lee decomposition of $r\cdot c$ is given by
(143) $\pi^{\mathsf{L}}(r\cdot c)=(\pi^{\mathsf{L}}_{0},\pi^{\mathsf{L}}_{{\sigma_{r}}(1)},\dots,\pi^{\mathsf{L}}_{\sigma_{r}(M-1)})=:\sigma(\pi^{\mathsf{L}}(r\cdot c)),$
where $\sigma_{r}(i)=k$ such that $rk\equiv i\pmod{p^{s}}$.
Lemma 39 ([3, Lemma 1]).
Given a Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, a unit element $r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$ and $\sigma_{r}$ a permutation on the set $\left\{{1,\ldots,M}\right\}$, we define $\sigma_{r}(\pi):=(\pi_{0},\pi_{{\sigma_{r}}(1)},\dots,\pi_{\sigma_{r}(M-1)})$. Then, for any linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$, it holds that $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\sigma(\pi)}^{\mathsf{L}}(\mathcal{C})$.
For any Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, we define the set
(144) $S(\pi):=\left\{{\sigma_{r}(\pi)\;|\;r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times},\ \sigma_{r}(i)=k\text{ with }kr\equiv i\pmod{p^{s}}}\right\}.$
Since $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})$ for every $\rho\in S(\pi)$, an additional constraint arises in the formulation of the LP problem for linear Lee-codes.
Let $B_{p,1,s,n}(n,d_{\mathsf{L}})$ denote the maximum number of codewords in a linear code of length $n$ and minimum Lee-distance $d_{\mathsf{L}}$ over $\mathbb{Z}/{p^{s}}\mathbb{Z}$.
Theorem 40 (Linear Programming Bound for Lee-metric Codes).
For a linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ of minimum Lee distance $d_{\mathsf{L}}$ it holds that
(145) $B_{p,1,s,n}(n,d_{\mathsf{L}})\leq\max\left(\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})\right),$
where the maximum is taken over all $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})$ subject to the following conditions:
  • $\mathcal{D}_{\pi^{(0)}}^{\mathsf{L}}(\mathcal{C})=1$;
  • $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})\geq 0$ for all $11\leq i\leq D-$ 1 and $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})=0$ for all $i\in I$;
  • $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\pi^{(j)}}^{\mathsf{L}}(\mathcal{C})$ for all $\pi^{(j)}\in S(\pi^{(i)})$;
  • and
    (146) $\sum_{i=1}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})K^{\mathsf{L}}_{\pi^{(j)}}(\pi^{(i)})\geq\begin{cases}-\binom{n}{\pi^{(i)}_{1},\ldots,\pi^{(i)}_{M}}2^{n-\pi^{(i)}_{0}-\pi^{(i)}_{M}}&\text{if }p=2,\\ -\binom{n}{\pi^{(i)}_{1},\ldots,\pi^{(i)}_{M}}2^{n-\pi^{(0)}_{i}}&\text{otherwise,}\end{cases}$
where $I=\{i\ \;|\;\ x\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}\text{ such that }\pi(x)=\pi^{(i)}\text{ and }\operatorname{wt}_{\mathsf{L}}(x)<d_{\mathsf{L}}\}$.
The last constraint is derived from the fact that
(147) $\mathcal{D}_{\pi^{(j)}}^{\mathsf{L}}(\mathcal{C}^{\perp})=\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})K^{\mathsf{L}}_{\pi^{(j)}}(\pi^{(i)})\geq 0$
and from computing the Krawtchouk coefficient of $\pi^{(0)}$ with respect to $\pi^{(j)}$.
However, note that this formulation of the LP problem still does not result in an efficient program due to the high number of variables involved.
5. Conclusion
In this paper we proved the existence of MacWilliams-type identities for any finite chain ring and any additive weight. In particular, this includes the homogeneous, the Lee and the subfield metric, where the classical MacWilliams identities fail. The main idea of the novel identities is to partition the ambient space into a finer partition, called Lee partition which is able to recover the weight enumerator of any additive weight. We then improved upon this result in the case of the homogeneous and the subfield metric, by presenting a coarser partition than the Lee partition, but finer than the weight partition, for which MacWilliams-type identities hold. The presented identities can be used to formulate LP bounds in each of the metrics.
In this work we laid the main focus on the Lee, homogeneous and subfield metric. Even though the MacWilliams-type identities for the Lee partition enumerator allows to derive the weight enumerator of any additive weight over any finite chain ring, for specific weights there are more natural choices on how to partition the code. Future work, therefore, includes the study of other weights, such as the Bachoc weight [7], Krotov weight [29, 38] or the overweight [17].
Acknowledgments
Jessica Bariffi was partly supported by the Federal Ministry of Education and Research of Germany in the program of "Souverän. Digital. Vernetzt." Joint project 6G-RIC, project identification number: 16KISK022. She is now funded by the European Union (DiDAX, 101115134). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Giulia Cavicchioni is member of the INdAM Research Group GNSAGA. Violetta Weger is supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 899987.